### v2

# An LLM-based Agent for Reliable Docker Environment Configuration

Abstract
Environment configuration is a critical yet time-consuming step in software development, especially when dealing with unfamiliar code repositories. While Large Language Models (LLMs) demonstrate the potential to accomplish software engineering tasks, existing methods for environment configuration often rely on manual efforts or fragile scripts, leading to inefficiencies and unreliable outcomes. We introduce Repo2Run, the first LLM-based agent designed to fully automate environment configuration and generate executable Dockerfiles for arbitrary Python repositories. We address two major challenges: (1) enabling the LLM agent to configure environments within isolated Docker containers, and (2) ensuring the successful configuration process is recorded and accurately transferred to a Dockerfile without error. To achieve this, we propose atomic configuration synthesis, featuring a dual-environment architecture (internal and external environment) with a rollback mechanism to prevent environment “pollution” from failed commands, guaranteeing atomic execution (execute fully or not at all) and a Dockerfile generator to transfer successful configuration steps into runnable Dockerfiles. We evaluate Repo2Run on our proposed benchmark of 420 recent Python repositories with unit tests, where it achieves an 86.0% success rate, outperforming the best baseline by 63.9%. Repo2Run is available at https://github.com/bytedance/Repo2Run.
1 Introduction
Figure 1: The illustration of a command executing failed and “polluting” the environment.
Figure 2: Examples of the failed command causing “polluted” environment.
Configuring the environment is typically the first step for developers to run and test the code. Meanwhile, Large Language Models (LLMs) are increasingly being integrated into tools such as chatbots and coding assistants, showcasing their potential to accomplish various software engineering tasks [32, 12, 4]. As a result, the research community explores how LLMs can be leveraged to assist with more complex real-world tasks in software development. Furthermore, LLM-based agents [13, 5, 1] are also increasingly being utilized in software engineering. With the advancement of LLM-based agents and their applications in automated programming, many techniques now require running tests in the runtime environment to acquire more comprehensive information. However, previous approaches often rely on manually configuring environments [14], and writing individual configuration scripts for each task, which is both time-consuming and labor-intensive.
We conduct an empirical study with eight professional developers from Internet companies, all with an average of seven years of development experience. The results indicate that when dealing with unfamiliar code repositories, five out of eight developers spend a great amount of time configuring the environment. Moreover, there is a strong willingness among them to use automated tools to alleviate the burden of environment setup. Details of the empirical study are presented in Appendix A.
In response to the growing complexity and time required for manual environment configuration, containerization tools like Docker have become a good solution. Docker allows developers to create isolated and consistent environments across different systems, thus eliminating the common “it works on my machine” problem [23]. Dockerfiles are commonly used to configure Docker environments, which include two parts: the base image and the configuration process. The base image is a pre-packaged environment (e.g., python:3.10) that provides a standardized, isolated environment for development without the need for local deployment. The configuration process involves additional steps performed on top of the base image.
Inspired by the idea of the Docker container, we design our LLM agent for environment configuration, Repo2Run, with a fundamental concept: utilize an LLM agent to configure the environment within an isolated Docker container. If the LLM successfully configures the environment, we only need to record the configuration steps and transfer them into a Dockerfile. During the environment configuration process, we mainly face the following two major challenges:
Ensure the LLM agent configures the environment within a Docker container. One major challenge is ensuring that the LLM agent can accurately configure the environment inside the Docker container. This involves selecting an appropriate base image, installing dependencies, resolving conflicts, and running tests, etc. If we do not provide the LLM agent with effective tools, the LLM agent will find it difficult to handle the various complex issues encountered during configuration.
Ensure the successful configuration process is recorded and accurately transferred to a runnable Dockerfile. Another major challenge is maintaining an accurate and comprehensive record of the environment configuration process. As illustrated in Figure 1, a failed executed command “commandA” may cause irreversible ”pollution”. Although “commandA” fails to run, it still causes ”pollution” to the original environment, potentially leading to unexpected changes to packages, files, or directories in the system. To accurately reproduce such system changes, we need to add the “RUN commandA” statement in the Dockerfile. However, this will cause the Dockerfile to build fail.
Figure 2 shows two practical examples. As shown in Figure 2 (a), when we run “pip install tensorflow-gpu”, the command fails to execute because the package tensorflow-gpu has been deprecated. However, it still introduces the “python-version” package which is not present in the original environment. In Figure 2 (b), when we run the “rm -rf /path/to/logs” command to delete the directory, some files fail to be deleted due to a lack of permission. However, the files that have already been deleted (e.g., log1.txt and log2.txt) are not restored. Such “pollution” caused by failed commands may make the environment enter an uncertain state. This state cannot be reached by directly adding commands to the Dockerfile, as doing so would cause the Dockerfile build process to fail.
To address the aforementioned two challenges, we propose Repo2Run based on atomic configuration synthesis, which consists of a dual-environment architecture and a Dockerfile generator. To the best of our knowledge, this is the first work that leverages the LLM agent for automated Dockerfile generation and environment configuration in a completely automated manner.
The dual-environment architecture comprises both an internal and an external environment. The internal environment serves as a Docker container sandbox, allowing the LLM agent to execute various commands. The external environment assists in configuring the environment in the internal environment. It facilitates the action-observation interactions of the LLM agent and supports operations such as changing the base image of the internal environment. In addition, we design a rollback mechanism that performs a rollback when a command fails, preventing the system from entering an uncertain state. This ensures that we can execute commands “atomically”, meaning each command either executes fully or not at all.
The Dockerfile generator designs a set of rules to transfer the successfully executed commands in the internal environment into statements in the Dockerfile without errors. Combined with the rollback mechanism, we finally ensure that all generated Dockerfiles are successfully built.
In summary, the contributions of this paper are as follows:
$\bullet$ We propose Repo2Run, the first LLM-based agent designed to automate environment configuration and generate Dockerfiles for arbitrary Python repositories, enabling the execution of tests within these repositories.
$\bullet$ We propose atomic configuration synthesis, which includes a dual-environment architecture and a Dockerfile generator, ensuring that all Dockerfiles generated by Repo2Run are runnable.
$\bullet$ We create a benchmark to evaluate the performance of Repo2Run by crawling 420 popular Python repositories with unit tests from GitHub released within 2024. Repo2Run successfully configures the environment for 361 of these repositories, achieving a success rate of 86.0%, which is 63.9% higher than the best baseline.
2 Environment configuration
In this part, we formulate the task of the environment configuration. It involves identifying an appropriate base image ($B$) and a configuration process ($\Gamma$). The objective is to ensure that the final system state ($S_{f}$) successfully runs all the unit tests in the repository.
2.1 State Transition
Environment State ($S$): The environmental state ($S$) represents the current state of the computer system, which encompasses all variables, files, cache, etc. Any state $S\in\mathcal{S}$, where ($\mathcal{S}$) denotes the set of all possible environment states.
Command ($C$): The command $C$ represents an individual instruction or action that can be executed in the environment through interfaces such as bash, thus directly changing the system state.
State Transition Function ($\delta$): The state transition function ($\delta$) defines the process through which the system transitions from one state to another state upon execution of a given command, i.e.,
$\delta:\mathcal{S}\times\mathcal{C}\rightarrow\mathcal{S},\quad\delta(S,C)=S^{\prime}$ (1)
Here, $S^{\prime}$ is the new state after execution of command $C$.
2.2 Base image
Empty initial state ($S_{\emptyset}$): The empty initial state ($S_{\emptyset}$) represents a completely unconfigured, bare operating system or a purely hypothetical state without applications and configurations. This state is used for theoretical purposes to define the starting point of a system’s configuration. $S_{\emptyset}\in\mathcal{S}$.
Base image ($B$): From the empty initial state ($S_{\emptyset}$), we can construct the base image $B$ by applying a series of actions. The base image $B$ is essentially the final state achieved by applying a sequence of instructions starting from $S_{\emptyset}$:
$B=\delta^{m}(S_{\emptyset},{C^{B}_{1},C^{B}_{2},\ldots,C^{B}_{m}})$ (2)
Here, ${C^{B}_{1},C^{B}_{2},\ldots,C^{B}_{m}}$ are the individual commands that constitute the base image.
2.3 Configuration process and target state
Configuration Process ($\Gamma$): The configuration process, denoted as $\Gamma$, involves transforming the environment starting from the base image $B$ by applying a series of actions. The result of this process is a state $S_{f}$, which we refer to as the resultant state. Formally, the configuration process can be represented as:
$S_{f}=\delta^{n}\left(B,{C_{1},C_{2},\ldots,C_{n}}\right)$ (3)
Following this transformation, it is crucial to verify whether $S_{f}$ meets all the required conditions to be considered the target state. This verification is performed using the state verification function.
State verification ($\epsilon$): The state verification function ($\epsilon$) determines whether the resultant state ($S_{f}$) successfully runs all tests within the repository after executing the configuration process ($\Gamma$). It maps a given state ($S$) to a binary value indicating whether all tests could be run in the resultant configuration: $\epsilon(S)$ returns 0 if all tests are successfully run, and 1 otherwise. If $\epsilon(S_{f})=0$, it indicates that the configuration process has finished and the resultant state $S_{f}$ is considered the target state. Otherwise, it means that $S_{f}$ does not meet the necessary conditions.
3 Repo2Run
Figure 3: An example process of Repo2Run, which illustrates two main parts: Event Stream and Repo2Run Workflow. The Event Stream tracks the entire action-observation process, where green boxes represent system prompts, yellow boxes represent responses from the LLM agent including the actions, and blue boxes represent observations from the Runtime Environment. The shaded actions indicate configurations abandoned after base image change. In the Event Stream, the blue text indicates that the command starts running, the green text indicates that the command runs successfully, and the red text indicates that the command fails. The Repo2Run Workflow consists of the internal environment and the external environment. The internal environment serves as the actual configuration Docker-based sandbox, which builds an actual testing runtime environment. The external environment executes the action-observation process and assists in the configuration process within the internal environment.
In this part, we introduce the design of Repo2Run and atomic configuration synthesis. As shown in Repo2Run Workflow in Figure 3, the dual-environment architecture consists of an external environment and an internal environment. The internal environment is a Docker container where the actual environment configuration is executed, while the external environment assists the internal environment in setting up the configuration. After successfully configuring the environment, the Dockerfile generator will create the Dockerfile.
Figure 4: An example of the Dockerfile generator to transfer the commands into a runnable Dockerfile.
3.1 External environment
$\bullet$ Action-observation interaction: In each turn of the environment-agent interaction, the event history $E_{t}={(a_{1},o_{1}),(a_{2},o_{2}),\ldots,(a_{t-1},o_{t-1})}$ records the historical actions and observations up to time step $t$. Additionally, the sequence of successfully executed commands ${C_{1},C_{2},\ldots,C_{n}}$ helps remind the LLM agent of the current system state.
At time step $t$, given the event history $E_{t}$ and the sequence of successfully executed commands $C_{1},C_{2},\dots,C_{n}$, the LLM agent generates a response including an action $a_{t}$:
$a_{t}=\pi(E_{t},{C_{1},C_{2},\ldots,C_{n}})$ (4)
where $\pi$ is the function of the LLM agent’s decision. The environment state then transitions based on the action $a_{t}$:
$S_{t+1}=\delta(S_{t},a_{t})$ (5)
This iterative process ensures that the LLM agent remains informed of the system’s current state and makes decisions accordingly, facilitating effective multi-turn interaction within the environment.
$\bullet$ Rollback ($\rho$): As illustrated in Figure 1, commands $C_{fail}$ may lead to an uncertain state ($S_{u}$) if not successfully executed. When a command executes failed (i.e., the command returns a non-zero return code), the system transitions into an uncertain state as defined by:
$S_{u}=\delta(S_{k-1},C_{fail})$ (6)
Here, $C_{fail}$ is the failed executed command. If a command fails, the system will transition to the uncertain state $S_{u}$. To avoid such uncertainty, we introduce rollback ($\rho$). In the event of a failure at any step, the rollback mechanism restores the environment to the last certain state:
$\rho:\mathcal{S}\times\mathcal{C}\rightarrow\mathcal{S},\quad\rho(S_{u},{C_{1},C_{2},\ldots,C_{k}})=S_{k-1}$ (7)
Specifically, each time an instruction ($C$) is executed, we utilize the “docker commit” command to take a snapshot of the current state. Once the instruction is completed, we check its return code. If the return code is not 0, it indicates a failure in executing the instruction, and we replace the image with the most recently committed one. Besides, some instructions, such as “cat”, generally do not change the state of the environment, and thus, we do not apply the rollback mechanism to them. The complete list of these actions is shown in Appendix B.
$\bullet$ Base image change: If the LLM agent identifies that the currently selected base image is incorrect or unsuitable during the environment configuration process, it can reselect the base image. Once a new image is chosen, all previous configurations become invalid, and the executed commands are cleared, necessitating a restart of the configuration process. For instance, as shown in Figure 3, when the LLM agent discovers that the “StrEnum” class cannot be imported through testing, which is only available in the Python standard library starting from version 3.11, it decides to reselect the base image to python:3.11.
$\bullet$ Result processor: In each turn of the interaction, executing commands may generate extensive output (e.g., logs), which can greatly increase the burden of the LLM agent. Therefore, whenever such information is added to the event history, we truncate excessively long text, retaining only the initial and final sections up to a specific length.
$\bullet$ Dockerfile generator: If the LLM agent successfully runs all tests in the repository in the internal environment, the configuration process needs to be recorded in a Dockerfile. Figure 4 illustrates the Dockerfile generator’s operation. 4 (a) lists the internal commands actually executed in the internal environment, their return codes (e.g., 0 indicates success, otherwise failure), and the current directory (dir). 4 (b) shows the process of generating the Dockerfile.
The principles for the conversion process include the following steps, scanning all internal commands sequentially, and the line numbers mentioned below match the line number of Figure 4 (a):
  1. 1.
    By default, use python:3.10 as the base image.
  2. 2.
    For commands that run successfully (i.e., have a return code of 0), prepend the command with “RUN” to form a Dockerfile statement (e.g., line 3). Commands that fail (i.e., have a return code other than 0, such as line 4) are rolled back and not included in the Dockerfile. An exception is some commands typically do not change the current state (such as “cat” on line 2). These commands are not added to the Dockerfile. We show the details in Appendix B.
  3. 3.
    If dir is not the root directory, use cd to change to that directory before running the command, as each Dockerfile statement runs in its own directory session.
  4. 4.
    If a base image change statement is encountered (e.g., line 5), discard all previous configurations and switch to the new base image.
  5. 5.
    If a code editing command is encountered (e.g., line 9), copy the patch and editing script into the Docker container before executing the editing command.
  6. 6.
    If an export statement for adding environment variables is encountered (e.g., line 13), convert it to a persistent ENV statement.
  7. 7.
    After all statements have been scanned, check for dependency installation commands and replace any unspecified versions (e.g., “B>=1.0,<2.0” in line 11) with the actual versions downloaded in the Docker container (e.g., “B==1.5.1” in the final Dockerfile of Figure 4 (b)).
3.2 Internal environment
As shown in the Repo2Run workflow of Figure 3, the internal environment consists of a Docker container with a base image $B$ and the code repository to be configured. According to the latest data [^1] https://w3techs.com/technologies/history_details/pl-python/3 from 2025, we select Python 3.10 as our default Docker base image due to its increasing usage and wide application among all Python 3 versions. Additionally, we categorize the executable actions within the internal environment into five types: environment monitoring, dependency installation, bash commands, code editing, and test running.
$\bullet$ Environment monitoring: It serves as the eyes of the LLM agent in the internal environment, allowing it to observe the current state of the environment. These commands usually do not change the state of the environment. For instance, commands like “ls” and “cat” are employed to inspect directories and files. Additionally, commands like “pip list” and “pipdeptree [^2] https://github.com/tox-dev/pipdeptree are utilized to get the installed version of third-party libraries in the current environment. For example, in the fourth part of Event Stream in Figure 3, the LLM agent calls the “find” command to search for requirements.txt within the environment to determine which dependencies need to be installed.
$\bullet$ Dependency installation: It is primarily used to install the third-party libraries necessary for running tests, including libraries managed by both “pip” and “apt-get”. To avoid dependency conflicts that may arise when downloading multiple packages, we implement a dependency management system in the external environment. Initially, the third-party libraries to be downloaded are added to a waiting list. If there are different constraints for the same library (e.g., “A>=>=> =1.0” and “A$<$1.0”), they will be moved to a conflict list, and the LLM agent will determine the appropriate version to download.
We design a set of tools for the LLM agent to manage the installations. For instance, as shown in Figure 3, the action “waitinglist addfile” adds all elements from requirements.txt to the waiting list, while executing the “download” command, third-party libraries are taken one by one from the waiting list and installed in the internal environment. The installation commands (e.g., “pip install” and “apt-get install”) are executed. If the installation succeeds, the state of the environment is updated; if it fails, a rollback is performed. We show all our designed tools in Appendix C.
$\bullet$ Test running: It detects whether the current environment is successfully configured by running “pytest” unit tests. If the state ($S_{f}$) of the current environment passes the tests, it indicates that the LLM agent has successfully configured the environment within the Docker container, and the internal environment’s process concludes. Conversely, if some tests fail, the error information is communicated back to the LLM agent, allowing it to perform further configurations and adjustments based on the errors.
$\bullet$ Code editing: It allows the LLM agent to modify the code within the internal environment, including codes inside and outside the repository. Besides, to prevent the LLM agent from spuriously running the tests by directly modifying or deleting the test files, we do not permit the LLM agent to modify or delete the original test files within the repository. For instance, Figure 3 demonstrates code editing to correct a syntax error resulting from the improper use of double quotes in a Python f-string.
$\bullet$ Bash commands: Aside from the commands above, we also enable the LLM agent to directly invoke bash commands to interact with the internal environment. As shown in Figure 3, the LLM agent uses the command “export PYTHONPATH=/repo/src”.
4 Experiment
We evaluate the effectiveness of Repo2Run on 420 Python code repositories. As the popular option, we select gpt-4o-2024-05-13 for subsequent experiments, with the temperature uniformly set to 0.2.
4.1 Benchmark
To the best of our knowledge, there is no prior work similar to Repo2Run that constructs an environment for arbitrary Python repositories. To validate the capability of Repo2Run in environment configuration, we create our new benchmark consisting of selected Python repositories from GitHub based on the following criteria:
$\bullet$ Creation date: To minimize the impact of data contamination, we carefully selected repositories created in 2024, ensuring they are not part of the training data for mainstream large language models.
$\bullet$ Star count: To ensure the quality of the repositories, we only select those with a star count greater than 100.
$\bullet$ Test directory: Pytest is a leading Python testing framework that is compatible with many other Python testing tools, such as unittest. It identifies and runs test files and test functions that are named with a “test_” prefix or a “_test” suffix. Python repositories usually place their test files in the “test” or “tests” directory. To effectively filter repositories with unit tests, we keep repositories with the “test” or “tests” directory in their root directory to filter the repositories that most probably have unit tests.
Through the selection process, we crawl all 449 repositories that meet all the above requirements [^3] The crawling was performed in December 2024.. Then, we filter 420 repositories containing at least one unit test. These repositories constitute our benchmark for subsequent experiments. The details of the benchmark are presented in Appendix D.
4.2 Evaluation metrics
$\bullet$ Dockerfile Generation Success Rate (DGSR): It indicates the percentage of attempts where the method successfully generates a runnable Dockerfile. To be considered successful, the generated Dockerfile must be able to build without errors. If the Dockerfile for a code repository successfully builds, it is regarded as a successful generation. Generating runnable Dockerfile is fundamental for successfully configuring the environment.
$\bullet$ Environment Configuration Success Rate (ECSR): It represents the percentage of attempts where the method successfully configures environments. For a successful configuration, the generated Dockerfile must not only build successfully but also allow tests to run by “pytest” in the Docker container. We are only concerned with whether tests can be executed, regardless of whether they pass or fail, as outcomes of tests may inherently vary within the repository.
Table 1: The results of different baselines.
Metric DGSR ECSR
pipreqs 29.8% (125) 6.0% (25)
LLM generator 47.6% (200) 22.1% (93)
SWE-agent 26.9% (113) 9.0% (38)
Repo2Run 100% (420) 86.0% (361)
4.3 Baselines
$\bullet$ pipreqs [^4] https://github.com/bndr/pipreqs: It is an automated tool that generates a “requirements.txt” file by analyzing the import statements in the Python scripts and identifying the necessary dependencies without LLM. Using the requirements.txt file generated by pipreqs, we create a Dockerfile. The detail is provided in Appendix E.
$\bullet$ LLM generator: The “README” file in a code repository usually contains environment configuration instructions. Therefore, we directly drive the LLM to read the “README” file and generate an executable Dockerfile accordingly.
$\bullet$ SWE-agent [26]: SWE-agent establishes a custom agent-computer interface (ACI) that facilitates the LLM agent’s interaction with the repository environment by allowing actions such as reading files, editing files, and executing bash commands. Initially intended as an LLM agent for bug fixing, we preserve its framework and default settings, adjust its prompts, and use it as a baseline.
4.4 Experimental Results
The results of different baselines are presented in Table 1. We observe that Repo2Run consistently outperforms other baselines on both DGSR and ECSR. Repo2Run ultimately completed environment configuration for 361 code repositories, achieving an ECSR of 86.0%. It is 63.9% higher than the highest rate achieved by other methods, demonstrating great advantages. Due to the design of atomic configuration synthesis, Repo2Run successfully generates Dockerfiles that can be built successfully for all 420 code repositories, which other tools cannot guarantee.
For pipreqs, the main failures come from two reasons. First, generating the requirements.txt fails when there are issues within the repository, such as encoding errors or syntax errors in the files. This happens in 30 repositories (7.1%). Second, even when requirements.txt is generated, it might not download properly due to package version conflicts. This occurs in 265 repositories (63.1%). Besides, both the LLM generator and SWE-agent fail to ensure that the generated Dockerfile can be successfully built due to the lack of an ensuring mechanism. Surprisingly, the ability of SWE-agent, a general agent framework, to generate Dockerfiles is even weaker than simply letting the LLM read the “README” file. This indicates that a general agent framework cannot guarantee the generation of runnable Dockerfiles. Ensuring mechanisms like atomic configuration synthesis are necessary to effectively use the interactive information from the agent to generate runnable Dockerfiles.
Table 2: The results of ablation experiments.
Metric DGSR ECSR
w/o dual-environment 92.4% (388) 41.7% (175)
w/o Dockerfile generator 19.5% (82) 13.8% (58)
Repo2Run 100% (420) 86.0% (361)
4.5 Ablabtion of Repo2Run
To investigate the impacts of the dual-environment architecture and Dockerfile generator separately, as two parts of atomic configuration synthesis, we separately remove each component of them. For the experiment without the dual-environment architecture, we retain only the internal environment’s bash commands as the most basic interface and remove all other tools. For the experiment without the Dockerfile generator, we directly instruct the LLM to generate a runnable Dockerfile.
Experimental result of the ablation study is shown in Table 2. We observe that removing the dual-environment architecture and retaining only bash commands results in a 7.6% decrease in DGSR. The main reason for this drop is the removal of rollback and other designs, making the system more prone to entering uncertain states and subsequently failing to reproduce. In addition, ECSR shows a 44.3% decrease, primarily because the simplification of design makes it more difficult for the LLM agent to complete configuration in the internal environment. Besides, removing the Dockerfile generator directly leads to an 80.5% drop in DGSR. This indicates that having the LLM directly generate Dockerfiles is unlikely to fully follow the event history, resulting in Dockerfiles that fail to build successfully. This also directly causes a sharp decline in ECSR.
It is also observed that Repo2Run without the Dockerfile generator is outperformed by the LLM generator. This is because the LLM generator leverages the “README” file, which provides a clear, simple and high-level overview of the environment configuration, allowing for more accurate Dockerfile generation. In contrast, the event history-based approach lacks this context, making it harder for the LLM to fully understand the configuration goals. However, the Dockerfile generator effectively utilizes the detailed event history, highlighting the complementary roles of both components of Repo2Run in generating a reliable Dockerfile.
5 Discussion
5.1 Tool usage frequency
Figure 5: Tool usage frequency of the success configuration.
As shown in Figure 5, we analyze the invocation times of various action types in 361 successfully configured projects, including the five actions within the internal environment and the action of changing the base image. Bash commands are the most frequently used action, as they encompass the majority of instructions. Additionally, we observe that the LLM agent tends to call dependency installation quite frequently, averaging about 18 times per configuration, which means roughly 18 dependencies are installed per configuration on average. Moreover, the LLM agent calls test running approximately 3.5 times per configuration on average, which typically helps the agent better identify issues. Among successful configurations, we see 48 instances of changing the base image, accounting for 13.3%  of success cases, indicating that the initial selection of the base image is often incorrect and requires subsequent adjustments.
5.2 Time consumption
Figure 6: The time distribution of each successful configuration.
Figure 6 shows the distribution of time spent successfully configuring 361 code repositories. The average time for successful configuration using Repo2Run is 29.03 minutes. 111 (30.7%) of the repositories are successfully configured in less than 10 minutes. Additionally, our empirical study through sampling indicates an average manual configuration time of 21.33 minutes (Appendix A). Considering network differences and randomness, Repo2Run achieves a time consumption comparable to manual configuration. Additionally, for complex issues, Repo2Run shows greater advantages over manual configuration. Repo2Run successfully configures all the cases that were manually successful in our empirical study. Moreover, the cases where Repo2Run fails are also not successfully configured manually.
5.3 Case study
For the code repositories that fail to configure, we manually inspect the reasons for failures and find that most are due to issues within the repositories themselves. Figure 7 illustrates a “module not found” error when configuring the repository “jialuechen/deepfolio”. In this case, the issue arises from the absence of updating the unit tests. The test file “test_stats.py” attempts to import modules from “deepfolio.stats”, but “stats” does not exist in “deepfolio”. Consequently, no matter how the LLM agent operates, it cannot directly run this test. This highlights the importance for developers to continuously update existing unit tests as the code repository evolves.
6 Related Work
6.1 Environment configuration and Dockerfile generation
Oss-Fuzz-Gen [17] relies on predefined build instructions (such as “./bootstrap.sh”, “./configure”, “make”) to build projects for fuzzing but lacks of flexibility for diverse projects when specified files are absent.
Existing solutions that help developers write Dockerfiles broadly fall into three categories: (1) Template-based generators that create Dockerfiles based on project context [2, 18], (2) Task-specific tools such as DockerizeMe which supports environment dependency inference for Python projects  [11] and DockerGen  [27] for dependency recommendations based on knowledge graphs built from existing Dockerfiles, and (3) Code completion tools including GitHub Copilot [7] and HumpBack [8] that generate suggestions for developers while writing Dockerfiles. The use of deep learning models to generate Dockerfiles based on natural language specifications of software requirements is also investigated [21]. While these approaches provide valuable assistance, they either require significant manual input from developers or are limited to specific use cases.
6.2 LLM-based agent
LLM-based agents typically consist of four key components: planning, memory, perception, and action [25]. Planning is crucial for agent systems, as it schedules agents to ensure a smooth process. LLM-based agents employ various planning strategies, including single [3] or multiple planners [15], single [29] or multi-turn planning [33], and single [9] or multi-path planning [31]. The memory component in LLM-based agents stores historical data to support coherent reasoning and complex tasks. Implementations vary in terms of memory duration (short-term [6] or long-term memory [28]), memory ownership (specific [20] or shared memory [10]). For perception, LLM-based agents primarily utilize textual input [19, 28] (natural language and programming language) and visual input [24, 22] (images and diagrams) to perceive and process information. To extend capabilities beyond interactive dialogue, the action component employs various external tools [30, 16], such as searching tools, file operations, and GUI operations. Nowadays, LLM-based agents have demonstrated superior performance compared to standalone LLMs in various software engineering tasks [13, 5, 1]. However, no LLM-based agents are specifically designed for environment configuration currently. To fill this gap, this paper employs a novel approach for automated coding environment configuration and Dockerfile generation.
Figure 7: An example of “module not found” error due to the absence of updating the unit tests in the repository.
7 Conclusion
In this paper, we propose Repo2Run, the first LLM-based agent for automated coding environment configuration and Dockerfile generation for Python repositories based on atomic configuration synthesis. With a dual-environment architecture and a Dockerfile generator, Repo2Run is able to select and change the base image, manage and install dependencies based on action observation and rollback mechanism, and utilize bash commands and existing test suites. Our evaluation of 420 popular Python repositories hosted on GitHub demonstrates the effectiveness of Repo2Run with an 86.0% success rate.
Impact Statement
Repo2Run demonstrates the potential and impact that LLM-based agents can have towards automating coding environment configuration and docker file generation. We are hopeful that Repo2Run may inspire more works that advance broader aspects of automated software engineering. There might be more potential societal consequences of our work, none of which which we feel must be specifically highlighted here.
Appendix
Appendix A Manual experiment
A.1 Settings
To ensure a fair comparison among participants, we conducted training and demonstrated examples before the experiment to ensure everyone understood the procedure. Additionally, to minimize discrepancies in time consumption due to network factors, all participants conducted the experiment in the same network environment.
A.2 Survey
We selected eight technical staff from internet companies to participate in the experiment and conducted a survey regarding their backgrounds prior to the experiment. Their development experience ranges from 4 to 11 years, with an average of 7 years in software development and 3.8 years in Python development. Seven participants have experience in complex development projects, while one has experience in multiple small-scale projects.
Regarding environment configuration, three participants indicated that they spend a significant amount of time configuring the environment when faced with an unfamiliar code repository; another three stated that, although they spend a long time, it is generally manageable; two participants reported spending minimal time.
In terms of successful environment configuration in their regular work, three participants mentioned they only fail with extremely complex environments, while five indicated they can build environments for most medium-scale repositories. As for their confidence in successfully configuring unfamiliar environments, six participants expressed that they are usually successful, and two said they are sometimes successful.
When it comes to the amount of time they are willing to wait to configure an unfamiliar repository, four participants are willing to wait for over 90 minutes, two are willing to wait 40-60 minutes, one is willing to wait 20-40 minutes, and one is only willing to wait 10-20 minutes.
Participants’ overall evaluation of configuring code running environments is diverse: one finds it very troublesome with many issues, four find it somewhat troublesome, two indicate moderate difficulty, and one finds it relatively simple.
Additionally, all participants expressed a high or very high willingness to use a tool that could automatically configure the environment for an unfamiliar repository.
A.3 Experiment guideline
We request all participants to conduct the experiment following the guidelines below:
1. Environment Setup
Configure the Docker environment. Verify the installation is successful: If installed correctly, you should be able to use the following command (python:3.10 is just an example; select the base image according to your requirements):
docker run -it python:3.10 bash
2. Overall Procedure
Our objective is to install the given package in a Docker container, configure its environment, and be able to run its internal tests. During the process, record the time developers spend configuring the environment and eventually save the logs using docker logs.
2.1 Review the Repository to be Configured (Optional)
Review the GitHub repository that needs to be configured.
2.2 Determine the Docker Base Image (Generally start with python:3.10)
Select the Docker base image based on the repository (all repositories use Python as the main language). Common base images include: Note, in this experiment, it is generally sufficient to use the official Python series images. The main concern is the version; if uncertain, you can start with newer versions like 3.10, 3.11, 3.12, or select the recommended Python version based on the repository’s README.
Python series (the number after python indicates the version): python:3.10, python:3.12, python:3.9, python:3.6, python:3.9… PyTorch series: pytorch/pytorch, pytorch/pytorch:1.9.0-cuda10.2-cudnn7-runtime… Anaconda series: continuumio/anaconda3… Note: If you find the selected version is incorrect later, you can exit and reselect.
2.3 Create and Enter the Container
Using the determined base image name (e.g., python:3.10), enter the container with the following command:
docker run -it --name mytest python:3.10 bash
Here, mytest is the container name, recorded for log export later. It can be freely named, just keep track of it to avoid losing it later. Note: If any issues arise here, check if Docker is correctly installed and if the image name is valid, and troubleshoot accordingly.
2.4 Install Relevant Tools
APT tool downloads:
apt-get update && apt-get install -y curl
Download pytest:
pip install pytest
2.5 Download the Repository
Select the repository to be configured and download its GitHub repository to a location (generally directly in the root directory):
git clone https://github.com/{full_name}.git
2.6 Configure the Environment
Now, use your skills to configure: First, enter the downloaded file directory, for example:
cd wddbfs
Switch to the specified branch SHA (refer to the corresponding SHA of the repository), for example:
git checkout 5c68aa
Our goal is to successfully run pytest (not necessarily to pass all tests, just to run them). A simple criterion is to successfully run:
pytest --collect-only -q
At this point, you can use your experience and information from the repository documentation and debugging error messages to configure. However, there are a few restrictions:
Do not directly edit the test files! (Files starting with test_ or ending with _test.py). Do not directly delete test files! Editing the original repository files is not recommended.
During this process, you may perform various operations, including but not limited to pip, apt-get, and other tool downloads, as well as searching online or using GPT for debugging help.
Additionally, if there are long download times requiring waiting, you may decide according to your situation whether to leave this running and do other things (just don’t forget about this task).
2.7 Completion and Logging
A task can conclude in two scenarios:
Scenario One: If “pytest --collect-only -q” runs without issues, you can then execute pytest. If pytest completes successfully, the task is done. Scenario Two: If you feel the package is extremely difficult to configure, for example exceeding your patience threshold (refer to your usual development habits), you may also terminate.
Once finished, input exit to exit. Make sure to save your output logs with the following command (replace container_name with the container name you recorded earlier, if you forget, you can use docker ps to check):
docker logs container_name -t > wddbfs.log
2.8 Fill Out the Form and Record Information
You need to fill out the form according to your feelings.
Question Description
Is it successful? Yes or No - whether the task successfully passed “pytest --collect-only -q” without errors and eventually ran “pytest”.
Final base image used E.g., python:3.10
Reason for failure Summarize the main reasons for failure (if failed), including: Long download time Difficulty handling repository dependencies Unresolvable bug Errors in repository tests Lengthy test durations Other reasons (please specify)
Waiting time Approximate value, indicating the time spent waiting for dependencies to download. Does not include time spent on decision-making and research. Provide an approximate range: ¡3 minutes 3-5 minutes 6-10 minutes 11-20 minutes 20-40 minutes 40-60 minutes 60-90 minutes 90+ minutes
Longest time-consuming process Describe the most time-consuming configuration process, such as: Downloading a specific dependency Resolving a specific error Incorrect Python version selection
Tolerance level Based on your subjective feeling and development experience. Rate the process (1-5): 1: Extremely unbearable 2: Somewhat unbearable, but manageable 3: Neutral, tolerable 4: Comfortable, no significant discomfort 5: Very comfortable, highly satisfactory configuration experience
Configuration difficulty Based on the complexity of the configuration process (ignoring time spent): Rate the process (1-5): 1: Very simple, completed with intuition and experience, no reference materials needed 2: Fairly simple, referred to basic materials (e.g., README), simple overall steps 3: Moderate difficulty, encountered some issues, but manageable 4: Difficult, required extensive debugging and configuration 5: Very difficult, needed numerous references and encountered unresolved or time-consuming issues
Materials referenced List the materials used for reference (e.g., README, internal configuration files, online searches, GPT). If none, mention “Directory”.
Biggest challenge during the process Describe the most troublesome aspect of the configuration process, such as long wait times, unclear error messages, dependency version conflicts, inability to find required software versions, etc.
Example 1:
Question Answer
Is it successful? Yes
Final base image used python:3.10
Reason for failure Successful
Waiting time 3-5 minutes
Longest time-consuming process Downloading poetry
Tolerance level 5
Configuration difficulty 1
Materials referenced Directory
Biggest challenge during the process Waiting for configuration and installation
Example 2:
Question Answer
Is it successful? No
Final base image used python:3.11
Reason for failure Unresolvable bug, provide bug image
Waiting time 40-60 minutes
Longest time-consuming process Resolving FileNotFoundError, ImportError, incorrect Python version selection
Tolerance level 2
Configuration difficulty 5
Materials referenced README, GPT, StackOverflow
Biggest challenge during the process Dependency version conflicts, long wait times
A.4 Repository assignment
We randomly assigned each participant four unique code repositories that were successfully configured by Repo2Run. Additionally, each participant was assigned two code repositories that Repo2Run failed to configure. To avoid chance occurrences, each failed repository was assigned to two different participants. Below is the list of selections:
Successfully configured:
[alexwlchan/safari-webarchiver, ManiMozaffar/aioclock, mixedbread-ai/batched, mobiusml/gemlite, circlemind-ai/fast-graphrag, knowsuchagency/promptic, mbodiai/embodied-agents, modelscope/agentscope, Adibvafa/CodonTransformer, kennethreitz/simplemind, lmstudio-ai/venvstacks, mlecauchois/micrograd-cuda, IST-DASLab/PanzaMail, MetaGLM/zhipuai-sdk-python-v4, openai/mle-bench, RealOrangeOne/django-tasks, basf/MolPipeline, dai-motoki/zoltraak, lucidrains/alphafold3-pytorch, mistralai/mistral-common, BMPixel/moffee, DataformerAI/dataformer, jahwag/ClaudeSync, volfpeter/htmy, Genentech/gReLU, OpenNLPLab/lightning-attention, paradigmxyz/spice, reagento/dishka, arcee-ai/fastmlx, KyanChen/RSMamba, neuralmagic/guidellm, simonw/files-to-prompt]
Failed to configure:
[zhuqinfeng1999/Samba, dongxuyue/Open-ReplaceAnything, LazyAGI/LazyLLM, jialuechen/deepfolio, KOSASIH/pi-nexus-autonomous-banking-network, AARG-FAN/Yolo_for_Wukong, plinder-org/plinder, expectedparrot/edsl]
A.5 Results
Based on the times shown in the logs, we calculated that the average configuration time for each repository is 21.33 minutes. Furthermore, none of the repositories that Repo2Run failed to configure were successfully configured manually. Additionally, in the manual experiment, five environments that were successfully configured by Repo2Run were not successfully configured, representing 15.6% of the total successfully configured assignments.
Appendix B Safe commands
If Repo2Run executes the following commands without using “$>$” or “>>much-greater-than$>$ >” for output redirection, they are regarded to be safe commands that typically do not affect the system. Therefore, rollback is not necessary, and they are not added to the generated Dockerfile.
[“cd”, “ls”, “cat”, “echo”, “pwd”, “whoami”, “who”, “date”, “cal”, “df”, “du”, “free”, “uname”, “uptime”, “w”, “ps”, “pgrep”, “top”, “dmesg”, “tail”, “head”, “grep”, “find”, “locate”, “which”, “file”, “stat”, “cmp”, “diff”, “xz”, “unxz”, “sort”, “wc”, “tr”, “cut”, “paste”, “tee”, “awk”, “env”, “printenv”, “hostname”, “ping”, “traceroute”, “ssh”]
Appendix C Repo2Run tools
Showing in Table 6, we design the following actions for Repo2Run to facilitate its invocation.
Table 6: Command list and their functions
Command Function
waitinglist add -p package_name Add item into waiting list. If no “version_constraints” are
[-v version_constraints] -t tool specified, the latest version will be downloaded by default.
waitinglist addfile file_path Add all entries from a file similar to requirements.txt format to
the waiting list. Format should be package_name [version_constraints].
waitinglist clear Clear all items in the waiting list.
conflictlist solve -v Resolve the conflict for the first element in the conflict list, and
[version_constraints] update the version constraints for the corresponding package_name and tool to version_constraints. If no “version_constraints” are specified, the latest version will be downloaded by default. The package_name and tool in the original waiting list must match one of the elements in the conflictlist. Here, the version_constraints are specified.
conflictlist solve -u Keep the original version constraint that exists in the waiting list, and discard the other version constraints with the same name and tool in the conflict list.
conflictlist clear Clear all items in the conflict list.
conflictlist show Show all items in the conflict list.
waitinglist show Show all items in the waiting list.
download Download all pending items in the waiting list at once, and the conflict list must be empty before executing.
runtest Check if the configured environment is correct using“pytest”.
poetryruntest Check if the configured environment is correct in the poetry environment. If you want to run tests in the poetry environment, run it.
runpipreqs Generate requirements_pipreqs.txt and pipreqs_output.txt and pipreqs_error.txt.
change_python_version Switching the Python version in the Docker container will forgo
python_version any installations made prior to the switch. The Python version number should be represented directly with numbers and dots, without any quotation marks.
clear_configuration Reset all the configuration to the initial setting of python:3.10.
Appendix D Benchmark
Table 7: Success status of each package in the benchmark.
full_name sha success full_name sha success
271374667/VideoFusion 9ba7b8 Yes 6abd/horus c1d093 Yes
a-r-r-o-w/cogvideox-factory 80d115 Yes a-s-g93/neo4j-runway 16b441 Yes
Aaditya-Prasad/consistency-policy eed0c4 No AARG-FAN/Yolo_for_Wukong 07f61a No
adamobeng/wddbfs 5c68aa Yes Adibvafa/CodonTransformer 2842ef Yes
AdityaNG/kan-gpt 0c6e4c Yes Admyral-Security/admyral de332e Yes
AgentOps-AI/AgentStack ff9c6a Yes aidatatools/ollama-benchmark c6a5fd Yes
AIR-Bench/AIR-Bench 4b27b8 Yes airbytehq/PyAirbyte 7e65ab Yes
airtai/fastagency 1ff503 Yes Akkudoktor-EOS/EOS fff685 Yes
alexmolas/microsearch 632ff2 Yes alexwlchan/safari-webarchiver 0e4974 Yes
AlibabaPAI/llumnix b319b2 Yes All-Hands-AI/OpenHands 246107 Yes
alvin-r/databonsai 3f2b7c Yes amchii/tg-signer 926dbb Yes
andrewyng/aisuite 763996 Yes andrewyng/translation-agent e0fc60 Yes
AnswerDotAI/byaldi 4583c0 Yes AnswerDotAI/rerankers ecd1f6 Yes
antgroup/agentUniverse ed8f55 Yes apapiu/transformer_latent_diffusion 84a75e Yes
apify/crawlee-python 267063 Yes apple/ToolSandbox 1a1dc8 Yes
apple/ml-cross-entropy 1f3ebd Yes apple/ml-mdm 9a5632 Yes
arcee-ai/fastmlx fd37bc Yes argmaxinc/whisperkittools 03898f Yes
arvindrajan92/DTrOCR a10aa0 Yes astramind-ai/Auralis c357a1 Yes
atonderski/neuro-ncap ecdcf2 Yes aurelio-labs/semantic-chunkers 04acc2 Yes
AuvaLab/itext2kg 941a1d Yes awslabs/agent-evaluation 3df695 Yes
Azure/co-op-translator a4709e Yes Azure-Samples/rag-postgres-openai-python 61bde7 Yes
bananaml/fructose 5f24ec Yes basf/MolPipeline 2f9bae Yes
basf/mamba-tabular af1ea0 Yes beatzxbt/mm-toolbox 728e35 Yes
bellingcat/ShadowFinder 578d5a Yes Benexl/FastAnime 677f46 Yes
betaacid/FastAPI-Reference-App 8caeca Yes bhavnicksm/chonkie 990493 Yes
bigcode-project/bigcodebench aa634d Yes Bl3f/yato 4906b0 Yes
Blealtan/efficient-kan 7b6ce1 Yes block/goose c497a5 Yes
BMPixel/moffee 0e643d Yes boheumd/MA-LMM ffe9fa Yes
bytewiz3/TravelGPT b19b43 Yes CausalLearning/ReAct 7d3665 No
cfahlgren1/observers d46fdb Yes chaidiscovery/chai-lab b6e7fa Yes
character-ai/prompt-poet 466432 Yes cheahjs/palworld-save-tools 7dc2c7 Yes
chernyadev/bigym 72d305 Yes chrschy/fact-finder ca57d1 Yes
circlemind-ai/fast-graphrag 447511 Yes cloudflare/cloudflare-python 228479 Yes
codefuse-ai/CodeFuse-muAgent e93924 No codeintegrity-ai/mutahunter f88922 Yes
codematrixer/hmdriver2 c0d075 Yes codeskyblue/tidevice3 d83c34 Yes
codeskyblue/uiautodev eb8577 Yes Codium-ai/AlphaCodium eb7577 Yes
Codium-ai/cover-agent 5c4b88 Yes COLA-Laboratory/TransOPT de8bf3 No
Comfy-Org/comfy-cli 7711db Yes CompEpigen/figeno 14b904 Yes
computer-agents/agent-studio d7f6cb Yes cosmic-cortex/mlfz 5bf8d2 Yes
cremebrule/digital-cousins 49400b Yes crewAIInc/crewAI-tools 873935 Yes
cvg/nerf-on-the-go 3659e7 Yes D-Star-AI/dsRAG 2d5431 Yes
D4Vinci/Scrapling 012820 Yes DAGWorks-Inc/burr 79137e Yes
dai-motoki/zoltraak 4dce44 Yes darrenburns/posting 94feab Yes
DataformerAI/dataformer 0cf88c Yes daxa-ai/pebblo e67b01 Yes
daya0576/beaverhabits c01257 Yes dbos-inc/dbos-transact-py d6c6ac Yes
deepsense-ai/db-ally 26033f Yes dendrite-systems/dendrite-python-sdk 27c9da Yes
denser-org/denser-retriever 76256e No dingo-actual/infini-transformer 08d0a1 Yes
discord/access 19e9b1 Yes dleemiller/WordLlama e38d47 Yes
dongxuyue/Open-ReplaceAnything 83f0ae No dottxt-ai/outlines-core 31ab9f Yes
dottxt-ai/prompts 3d2689 Yes dreadnode/rigging 82ac80 Yes
droid-dataset/droid_policy_learning 205ff6 Yes DS4SD/docling aee9c0 Yes
dynamiq-ai/dynamiq 6cca1c Yes eakmanrq/sqlframe 61fda5 Yes
EleutherAI/sae 0483b5 Yes emcf/thepipe 02e397 Yes
Emerging-AI/ENOVA b3661d Yes EnhancedJax/Bagels d72d7f Yes
enoch3712/ExtractThinker 4872a7 No epic-open-source/seismometer b3e812 Yes
epistoteles/TensorHue 1564fa Yes epogrebnyak/justpath 0aca51 Yes
erezsh/reladiff d8683b Yes etianen/logot 54e5ef Yes
expectedparrot/edsl aa7a2d No explosion/spacy-layout 64c6f4 Yes
facebookresearch/audioseal ea10f5 Yes facebookresearch/lightplane 34fe6c Yes
facebookresearch/spiritlm 52fb2f Yes FalkorDB/GraphRAG-SDK 250ebe Yes
Fanqi-Lin/Data-Scaling-Laws bd6941 No fastapi/fastapi-cli bc0840 Yes
fedirz/faster-whisper-server cbb6c9 Yes felafax/felafax 34a475 Yes
filipstrand/mflux 627398 Yes FlagOpen/FlagGems ca13b7 No
fmind/cookiecutter-mlops-package 00fef7 Yes foundation-model-stack/fms-fsdp 408c75 Yes
fpgmaas/cookiecutter-uv 90de47 Yes frdel/agent-zero 3cefa1 Yes
frostming/tetos 106ea5 Yes Fugaku-LLM/DeepSpeedFugaku 74753f No
gauge-sh/bridge 8b3430 Yes Genentech/gReLU efd308 Yes
genomoncology/FuzzTypes d96243 Yes getludic/ludic a6db96 Yes
getzep/graphiti 9f3dd5 Yes GigaxGames/gigax c3c209 Yes
gojasper/flash-diffusion 48e3bc Yes gomate-community/TrustRAG 1334c4 Yes
gomate-community/rageval 01e258 Yes google-deepmind/nanodo 10aefd Yes
google-deepmind/penzai fda6cd Yes google-deepmind/treescope dac18b Yes
google-research/timesfm 02bc2f Yes goombalab/hydra b6b9b7 Yes
gpustack/gpustack 4f0c67 Yes gregpr07/browser-use 5e545d Yes
groq/groq-python fa2e13 Yes gusye1234/nano-graphrag 18fa3a Yes
hailo-ai/hailo-rpi5-examples 82cfc8 Yes Haiyang-W/GiT ef2b64 No
HanaokaYuzu/Gemini-API e8a2d2 Yes HATTER-LONG/Verbiverse 82f988 Yes
hinthornw/trustcall eaaaad Yes HKUDS/HiGPT 2b0793 No
HKUDS/UrbanGPT be3515 No hngprojects/hng_boilerplate_python_fastapi_web bc9740 Yes
hpcaitech/Open-Sora 38de63 Yes hpcaitech/SwiftInfer 239fd3 No
hrnoh24/stream-vc faa629 Yes huchenlei/ComfyUI_omost 7ef00d Yes
huggingface/lerobot 4c41f6 Yes huggingface/lighteval 6ad727 Yes
HZAI-ZJNU/Mamba-YOLO ea97fc No IAAR-Shanghai/Grimoire 3fe89d Yes
ib-api-reloaded/ib_async 38cf54 Yes IBM/fastfit 396611 Yes
IBM/terratorch 16e5af Yes IEIT-Yuan/Yuan2.0-M32 b403a2 No
igorbenav/SQLModel-boilerplate 2ead04 Yes igorbenav/fastcrud dc831b Yes
igrek51/wat 0d6079 Yes illuin-tech/colpali e45c4c Yes
illuin-tech/vidore-benchmark 469665 Yes Indoxer/LKAN 16c48e No
Infini-AI-Lab/Sequoia 688079 No instanseg/instanseg 0df8b2 Yes
instructlab/instructlab c978b2 Yes Integuru-AI/Integuru 928e82 Yes
InternLM/InternEvo 5ad2eb Yes invariantlabs-ai/invariant 81547a Yes
IST-DASLab/PanzaMail b1807c Yes iterative/datachain b67d59 Yes
IvanDrokin/torch-conv-kan 7a0e83 Yes jahwag/ClaudeSync 000633 Yes
jgravelle/pocketgroq e995c4 Yes jhj0517/AdvancedLivePortrait-WebUI a7975c Yes
jialuechen/deepfolio 15d247 No jina-ai/late-chunking db558c Yes
jlowin/fastmcp baa300 Yes jmschrei/tangermeme a96897 Yes
jonbarron/camp_zipnerf 8e6d57 Yes JosephBARBIERDARNAL/pypalettes 826930 Yes
JoshuaC215/agent-service-toolkit c72f48 Yes jshuadvd/LongRoPE eb9aba Yes
jxnl/n-levels-of-rag 2ce110 Yes karpathy/minbpe 1acefe Yes
kennethreitz/simplemind 39b5a5 Yes kevinzakka/mink cf1a30 Yes
knowsuchagency/promptic a1930c Yes koaning/uvtrick 2d7f27 Yes
kohjingyu/search-agents 7c35ac No KOSASIH/pi-nexus-autonomous-banking-network 7fcff4 No
kotaro-kinoshita/yomitoku 71c85b Yes KruxAI/ragbuilder db3d3d No
kujirahand/tkeasygui-python b1f293 No KyanChen/RSMamba 3fa305 Yes
kyegomez/MultiModalMamba 58db40 No landing-ai/vision-agent 63eab8 Yes
langchain-ai/langchain-postgres 064e5b Yes lavague-ai/LaVague b3557f Yes
LazyAGI/LazyLLM e0dd38 No lenML/Speech-AI-Forge 0b31b2 Yes
leopiney/neuralnoise c0313f Yes lichao-sun/Mora 7a030e No
Lightning-AI/litdata 0a97de Yes lightonai/pylate 8de184 No
LilianHollard/LeYOLO 872841 Yes line/lighthouse ba9da7 Yes
LlmKira/fast-langdetect 5728ba Yes LMCache/LMCache 7d3443 Yes
lmstudio-ai/mlx-engine daeb7a Yes lmstudio-ai/venvstacks 235ce3 Yes
lucasdelimanogueira/PyNorch ed391e Yes LucasFaudman/apkscan 3b3e62 Yes
lucidrains/alphafold3-pytorch 49f7c9 Yes lucidrains/infini-transformer-pytorch 5774bb Yes
lucidrains/pi-zero-pytorch 8ad66f Yes lucidrains/titok-pytorch 2f9525 Yes
lucidrains/transfusion-pytorch 16f73e Yes MadcowD/ell 36ca5e Yes
ManiMozaffar/aioclock 3d196b Yes MaoXiaoYuZ/Long-Novel-GPT e952ac No
Marker-Inc-Korea/AutoRAG aa0bfb Yes martius-lab/hitchhiking-rotations 45b49f Yes
mbodiai/embodied-agents 8715f6 Yes McGill-NLP/weblinx 6f2014 Yes
McGill-NLP/webllama 696a7c Yes Menghuan1918/pdfdeal e08199 Yes
meta-llama/llama-stack-apps f14a73 Yes MetaGLM/zhipuai-sdk-python-v4 7ff4de Yes
metavoiceio/metavoice-src de3fa2 Yes microsoft/MInference 7a3e5a No
microsoft/TinyTroupe 9b8d4e Yes microsoft/Trace 826cf5 Yes
microsoft/aurora 8b1165 Yes microsoft/graphrag de1252 Yes
microsoft/semantic-link-labs 8e37ef Yes mikekelly/AgentK e9ec89 Yes
Mindinventory/MindSQL 3d0ff0 Yes MinishLab/model2vec 4e3fba Yes
miquido/draive 270f0c Yes mistralai/mistral-common 5cac5e Yes
mistralai/mistral-finetune 656df1 Yes mixedbread-ai/baguetter a6e915 Yes
mixedbread-ai/batched 1a1797 Yes mkjt2/lockbox 58430d Yes
mlecauchois/micrograd-cuda ab1ca0 Yes MLT-OSS/open-assistant-api 44eeaf Yes
mlx-graphs/mlx-graphs 4619d9 No mobiusml/gemlite 5ebcca Yes
ModelCloud/GPTQModel a5aefc No modelcontextprotocol/python-sdk aaf32b Yes
modelscope/MemoryScope 330b76 Yes modelscope/agentscope ceaf89 Yes
modern-python/that-depends 65e656 Yes muchdogesec/history4feed 614182 Yes
muditbhargava66/PyxLSTM f3c9bb Yes narwhals-dev/narwhals a2088f Yes
nasa-jpl/rosa 5471dc Yes neo4j/neo4j-graphrag-python 0ac06b Yes
neuralmagic/AutoFP8 e94461 Yes neuralmagic/guidellm ecf298 Yes
NewT123-WM/tnlearn 50ee75 Yes NexaAI/nexa-sdk 33f6ba No
nicobrenner/commandjobs 4c7264 Yes Nike-Inc/koheesio 9bd29e Yes
nlmatics/nlm-ingestor c72542 Yes NLPJCL/RAG-Retrieval d73057 No
nomic-ai/contrastors 496a05 No NousResearch/finetuning-subnet e2f5eb Yes
NUS-HPC-AI-Lab/VideoSys 6c92ae No NVIDIA/Megatron-Energon 28aa3b Yes
NVIDIA/NeMo-Skills 5591f3 Yes NVIDIA/kvpress 715f8a Yes
NVIDIA/logits-processor-zoo db179a Yes NVIDIA/nv-ingest eec9fa No
NVlabs/Sana 41dcbe Yes NVlabs/VILA ec7fb2 No
NVlabs/nvTorchCam cc27be Yes ogkalu2/comic-translate 1933d1 Yes
Open-Wine-Components/umu-launcher b0c0d4 Yes openai/mle-bench 51ec2b Yes
openai/swarm 9db581 Yes opendatalab/MinerU 391a99 Yes
openfoundry-ai/model_manager 34f9ff Yes opengeos/HyperCoast c1604c Yes
OpenInterpreter/aifs 3f74c6 Yes OpenNLPLab/lightning-attention d74395 Yes
openpsi-project/ReaLHF 62d9cd Yes openrecall/openrecall 225a27 Yes
OpenSPG/KAG 68b2c6 No orbital-materials/orb-models 251573 Yes
outspeed-ai/outspeed 049b40 Yes OwlAIProject/Owl 919226 Yes
PacktPublishing/LLM-Engineers-Handbook ec6717 Yes paradigmxyz/spice e962a9 Yes
patched-codes/patchwork c9b02b Yes paulrobello/parllama 421238 Yes
PeiJieSun/NESCL 365d20 Yes plinder-org/plinder 9658cc No
pomonam/kronfluence 884255 Yes PrefectHQ/ControlFlow f259fa Yes
PrimeIntellect-ai/OpenDiloco 71f5c2 Yes PrimeIntellect-ai/prime a974cf Yes
princeton-nlp/SWE-agent 8b3571 Yes proger/accelerated-scan db7145 Yes
pydantic/logfire 3d7924 Yes pymupdf/RAG b25718 No
pytorch-labs/LeanRL a416e6 Yes raphaelmansuy/code2prompt 3b377b No
RapidAI/RapidDoc 5e5fef Yes RapidAI/RapidLayout 8e9677 Yes
reagento/dishka 2ed985 Yes real-stanford/universal_manipulation_interface 298776 No
Realiserad/fish-ai f32c7f Yes RealOrangeOne/django-tasks e6d26c Yes
reidjs/text-scheduler 8bb7d6 Yes reka-ai/reka-vibe-eval 93ecd9 Yes
remigenet/TKAN 8a1de0 Yes rio-labs/rio eda40a Yes
robocasa/robocasa 27f992 Yes RobotecAI/rai d15910 No
robusta-dev/holmesgpt c4743a Yes royreznik/rexi f1dca8 Yes
run-llama/llama_deploy 47efff Yes run-llama/llama_extract 89438f Yes
run-llama/llama_parse f78186 Yes SamKhoze/ComfyUI-DeepFuze edd7fe No
ScrapeGraphAI/Scrapegraph-ai bae92b Yes seanchatmangpt/dspygen 69f305 No
serverless-ca/terraform-aws-ca 2da837 No ServerlessLLM/ServerlessLLM 8f1e6b Yes
ServiceNow/BrowserGym 12aa5e Yes ServiceNow/TapeAgents 3eca5c Yes
ServiceNow/WorkArena 0ab9cb No ShaShekhar/aaiela 4e8d6a No
shawntan/scattermoe 63b76a Yes ShoggothAI/motleycrew 19837e Yes
showlab/computer_use_ootb 419d9d Yes shun-liang/yt2doc 201ec2 Yes
siliconflow/BizyAir cdb3bb Yes simonw/files-to-prompt f9a4d8 Yes
simonw/llm-claude-3 c62bf2 Yes simonw/llm-cmd 74fb98 Yes
simonw/llm-jq beaada Yes simular-ai/Agent-S ca83be Yes
sirocco-ventures/raggenie 99dfe5 Yes souzatharsis/podcastfy 804a61 Yes
SpecterOps/cred1py 432f91 Yes StacklokLabs/promptwright 42f69b Yes
steinathan/reelsmaker 75369c Yes stephengpope/no-code-architects-toolkit ffc1a8 Yes
StonyBrookNLP/appworld bc9c47 Yes Storia-AI/sage f47fa4 No
superlinear-ai/raglite b02c5a Yes swarmzero/swarmzero 6fcd7a Yes
tahnok/colmi_r02_client 84d3a6 Yes taketwo/llm-ollama dd616e Yes
taobojlen/django-zeal 232987 Yes TencentARC/BrushNet 101dc3 No
TheAiSingularity/graphrag-local-ollama bcb98d Yes thousandbrainsproject/tbp.monty a39a26 Yes
THU-MIG/yolov10 6fbaf4 Yes thu-nics/MoA da034c No
tjmlabs/AgentRun 1997dd Yes tobiasfshr/map4d 0b8bcd Yes
Toloka/dbt-af e7f436 Yes TorchJD/torchjd 1eaafe Yes
tox-dev/tox-uv d7405a Yes TuragaLab/flybody 2e1088 Yes
turbo-llm/turbo-alignment 009574 Yes TY-Cheng/torchvinecopulib c3a477 Yes
ucbepic/docetl 00a761 Yes ultrasev/llmproxy 1a1100 Yes
uname-n/deltabase 5eafb9 Yes unifyai/unify ea2088 No
Vashkatsi/deply 6d6875 Yes VideoVerses/VideoTuna ffc6df Yes
vintasoftware/django-ai-assistant 5b26c7 Yes virattt/financial-datasets 985664 Yes
vllm-project/llm-compressor 606aab Yes volcengine/verl ed2eaf Yes
volfpeter/fasthx e850b9 Yes volfpeter/htmy 0322a3 Yes
vysakh0/dravid 25b03b Yes warmshao/FasterLivePortrait 6aa810 Yes
weareprestatech/hotpdf 55ab97 Yes web-arena-x/visualwebarena 89f5af No
Weixiang-Sun/Bora c08bb6 Yes whyhow-ai/knowledge-graph-studio c41043 Yes
whyhow-ai/rule-based-retrieval 91701f Yes whyhow-ai/whyhow 63a3c6 Yes
WongKinYiu/YOLO b96c8e Yes WU-CVGL/BAD-Gaussians bdd8b3 Yes
wy-z/container-vm 07d402 Yes xdit-project/xDiT a7bd51 Yes
xhluca/bm25s c4fef2 Yes yihong0618/klingCreator e567c6 Yes
yihong1120/Construction-Hazard-Detection f5e1ca Yes yinjunbo/IS-Fusion 86c882 No
YUCHEN005/GenTranslate 62e59d No YUCHEN005/RobustGER ad4e37 No
yurujaja/pangaea-bench e1d12e Yes ZenGuard-AI/fast-llm-security-guardrails 6a867c Yes
zeroasiccorp/logik ca4bb1 Yes zhuqinfeng1999/Samba 229687 No
zipnn/zipnn 007319 Yes zou-group/textgrad b2dc68 Yes
Appendix E pipreqs baseline settings
Figure 8 shows the template of a Dockerfile generated using “requirements_pipreqs.txt” created by pipreqs.
Figure 8: Dockerfile prompt using the “requirements_pipreqs.txt” generated by pipreqs.

### v3

@@
-An LLM-based Agent for Reliable Docker Environment Configuration
+Repo2Run: Automated Building Executable Environment for Code Repository at Scale
 Abstract
-Environment configuration is a critical yet time-consuming step in software development, especially when dealing with unfamiliar code repositories.
-While Large Language Models (LLMs) demonstrate the potential to accomplish software engineering tasks, existing methods for environment configuration often rely on manual efforts or fragile scripts, leading to inefficiencies and unreliable outcomes.
-We introduce Repo2Run, the first LLM-based agent designed to fully automate environment configuration and generate executable Dockerfiles for arbitrary Python repositories.
-We address two major challenges: (1) enabling the LLM agent to configure environments within isolated Docker containers, and (2) ensuring the successful configuration process is recorded and accurately transferred to a Dockerfile without error.
-To achieve this, we propose atomic configuration synthesis, featuring a dual-environment architecture (internal and external environment) with a rollback mechanism to prevent environment “pollution” from failed commands, guaranteeing atomic execution (execute fully or not at all) and a Dockerfile generator to transfer successful configuration steps into runnable Dockerfiles.
-We evaluate Repo2Run on our proposed benchmark of 420 recent Python repositories with unit tests, where it achieves an 86.0% success rate, outperforming the best baseline by 63.9%.
-Repo2Run is available at https://github.com/bytedance/Repo2Run.
+Scaling up executable code data is significant for improving language models’ software engineering capability.
+The intricate nature of the process makes it labor-intensive, time-consuming and expert-knowledge-dependent to build a large number of executable code repositories, limiting the scalability of existing work based on running tests.
+The primary bottleneck lies in the automated building of test environments for different repositories, which is an essential yet underexplored task.
+To mitigate the gap, we introduce Repo2Run, the first LLM-based agent aiming at automating the building of executable test environments for any repositories at scale.
+Specifically, given a code repository, Repo2Run iteratively builds the Docker image, runs unit tests based on the feedback of the building, and synthesizes the Dockerfile until the entire pipeline is executed successfully.
+The resulting Dockerfile can then be used to create Docker container environments for running code and tests.
+We created a benchmark containing 420 Python repositories with unit tests for evaluation.
+The results illustrate that Repo2Run achieves an 86.0% success rate, outperforming SWE-agent by 77.0%.
+The resources of Repo2Run are available at https://github.com/bytedance/Repo2Run.
 1 Introduction
+Figure 1: The pipeline of code repository mining and manual environment building.
+Developers manually write Dockerfiles through iterative steps including base environment selection, dependency installation, test running, error handling, validating the environment by running unit tests.
+Figure 2: Four error types (highlighted in different colors) that SWE-agent fails to resolve during environment building.
 Figure 3: The illustration of a command executing failed and “polluting” the environment.
-Figure 2: Examples of the failed command causing “polluted” environment.
-Configuring the environment is typically the first step for developers to run and test the code.
-Meanwhile, Large Language Models (LLMs) are increasingly being integrated into tools such as chatbots and coding assistants, showcasing their potential to accomplish various software engineering tasks [32, 12, 4].
-As a result, the research community explores how LLMs can be leveraged to assist with more complex real-world tasks in software development.
-Furthermore, LLM-based agents [13, 5, 1] are also increasingly being utilized in software engineering.
-With the advancement of LLM-based agents and their applications in automated programming, many techniques now require running tests in the runtime environment to acquire more comprehensive information.
-However, previous approaches often rely on manually configuring environments [14], and writing individual configuration scripts for each task, which is both time-consuming and labor-intensive.
-We conduct an empirical study with eight professional developers from Internet companies, all with an average of seven years of development experience.
-The results indicate that when dealing with unfamiliar code repositories, five out of eight developers spend a great amount of time configuring the environment.
-Moreover, there is a strong willingness among them to use automated tools to alleviate the burden of environment setup.
-Details of the empirical study are presented in Appendix A.
-In response to the growing complexity and time required for manual environment configuration, containerization tools like Docker have become a good solution.
-Docker allows developers to create isolated and consistent environments across different systems, thus eliminating the common “it works on my machine” problem [23].
-Dockerfiles are commonly used to configure Docker environments, which include two parts: the base image and the configuration process.
-The base image is a pre-packaged environment (e.g., python:3.10) that provides a standardized, isolated environment for development without the need for local deployment.
-The configuration process involves additional steps performed on top of the base image.
-Inspired by the idea of the Docker container, we design our LLM agent for environment configuration, Repo2Run, with a fundamental concept: utilize an LLM agent to configure the environment within an isolated Docker container.
-If the LLM successfully configures the environment, we only need to record the configuration steps and transfer them into a Dockerfile.
-During the environment configuration process, we mainly face the following two major challenges:
-Ensure the LLM agent configures the environment within a Docker container.
-One major challenge is ensuring that the LLM agent can accurately configure the environment inside the Docker container.
-This involves selecting an appropriate base image, installing dependencies, resolving conflicts, and running tests, etc.
-If we do not provide the LLM agent with effective tools, the LLM agent will find it difficult to handle the various complex issues encountered during configuration.
-Ensure the successful configuration process is recorded and accurately transferred to a runnable Dockerfile.
-Another major challenge is maintaining an accurate and comprehensive record of the environment configuration process.
-As illustrated in Figure 1, a failed executed command “commandA” may cause irreversible ”pollution”.
-Although “commandA” fails to run, it still causes ”pollution” to the original environment, potentially leading to unexpected changes to packages, files, or directories in the system.
-To accurately reproduce such system changes, we need to add the “RUN commandA” statement in the Dockerfile.
-However, this will cause the Dockerfile to build fail.
-Figure 2 shows two practical examples.
-As shown in Figure 2 (a), when we run “pip install tensorflow-gpu”, the command fails to execute because the package tensorflow-gpu has been deprecated.
-However, it still introduces the “python-version” package which is not present in the original environment.
-In Figure 2 (b), when we run the “rm -rf /path/to/logs” command to delete the directory, some files fail to be deleted due to a lack of permission.
-However, the files that have already been deleted (e.g., log1.txt and log2.txt) are not restored.
-Such “pollution” caused by failed commands may make the environment enter an uncertain state.
-This state cannot be reached by directly adding commands to the Dockerfile, as doing so would cause the Dockerfile build process to fail.
-To address the aforementioned two challenges, we propose Repo2Run based on atomic configuration synthesis, which consists of a dual-environment architecture and a Dockerfile generator.
-To the best of our knowledge, this is the first work that leverages the LLM agent for automated Dockerfile generation and environment configuration in a completely automated manner.
-The dual-environment architecture comprises both an internal and an external environment.
-The internal environment serves as a Docker container sandbox, allowing the LLM agent to execute various commands.
-The external environment assists in configuring the environment in the internal environment.
-It facilitates the action-observation interactions of the LLM agent and supports operations such as changing the base image of the internal environment.
-In addition, we design a rollback mechanism that performs a rollback when a command fails, preventing the system from entering an uncertain state.
-This ensures that we can execute commands “atomically”, meaning each command either executes fully or not at all.
-The Dockerfile generator designs a set of rules to transfer the successfully executed commands in the internal environment into statements in the Dockerfile without errors.
-Combined with the rollback mechanism, we finally ensure that all generated Dockerfiles are successfully built.
-In summary, the contributions of this paper are as follows:
-$\bullet$ We propose Repo2Run, the first LLM-based agent designed to automate environment configuration and generate Dockerfiles for arbitrary Python repositories, enabling the execution of tests within these repositories.
-$\bullet$ We propose atomic configuration synthesis, which includes a dual-environment architecture and a Dockerfile generator, ensuring that all Dockerfiles generated by Repo2Run are runnable.
-$\bullet$ We create a benchmark to evaluate the performance of Repo2Run by crawling 420 popular Python repositories with unit tests from GitHub released within 2024.
-Repo2Run successfully configures the environment for 361 of these repositories, achieving a success rate of 86.0%, which is 63.9% higher than the best baseline.
-2 Environment configuration
-In this part, we formulate the task of the environment configuration.
-It involves identifying an appropriate base image ($B$) and a configuration process ($\Gamma$).
-The objective is to ensure that the final system state ($S_{f}$) successfully runs all the unit tests in the repository.
+(a) Failed commands like “commandA” can irreversibly “pollute” the environment by altering packages, files, or directories, making subsequent builds unstable.
+(b) To reproduce the changes, such a failed command “RUN commandA” needs to be added to the Dockerfile.
+However, adding it will lead to building failure.
+Large language models (LLMs) have recently illustrated significant progress on solving software engineering issues [1], driving the advent of numerous coding LLM agents like MetaGPT [2], SWE-agent [3], OpenHands [4], Copilot [5], and Cursor [6].
+However, the lack of training environment remains a significant challenge to advancing LLMs in software engineering — As shown in Figure 1, while static code is readily accessible through the GitHub API, building executable environments for code testing and patch validation demands extensive manual effort even for skilled developers  [^1] We conducted a study to show how industry developers struggle in building environments.
+See Appendix E..
+This forces researchers to choose between (1) relying on unreliable static code metrics for rewarding (e.g., SWE-Fixer [7]) and SWE-RL [8], and (2) engaging developers to manually build limited repository environments (e.g., SWE-Gym [9] and SWE-Smith [10]) without scalability.
+Thus, an automated environment building infrastructure is urgently needed but remains underexplored.
+In this paper, we introduce Repo2Run, an LLM-based agent designed to build executable environments.
+The key idea is that if the agent can successfully navigate the process of building a repository’s dependencies and running its unit tests, its actions can be recorded and replayed to programmatically synthesize a runnable Dockerfile.
+This enables automated creation of isolated and consistent environments across different platforms via only one line command “docker build”, eliminating “it works on my machine” problem [11].
+Repo2Run addresses two key challenges in automating executable environment building:
+  1. 1.
+    Hard to explore a valid trajectory that can successfully build the environment.
+We found that without effective actions, the LLM agent would struggle to handle such a complex process.
+As shown in Figure 2, SWE-agent [3], a well-known LLM-based agent, often struggles with unresolved issues during environment building due to its lack of specialized actions.
+To address this challenge, Repo2Run uses an internal environment for building and an external environment for assistance, with nine actions designed to resolve these issues.
+  2. 2.
+    Failing to synthesize a runnable Dockerfile without execution failure.
+As shown in Figure 3, we observed that some incorrect commands in the trajectory can “pollute” the environment, leading to errors during synthesis.
+To tackle this challenge, we design an adaptive rollback action.
+When encountering failed commands like “commandA” that may cause “pollution” to the environment, Repo2Run restores the internal environment to its pre-execution state.
+This ensures the building process remains consistent and repeatable using the Dockerfile, avoiding unexpected errors.
+To demonstrate Repo2Run’s effectiveness, we first created a benchmark of 420 latest Python repositories with unit tests from GitHub in 2024, as existing datasets include only a very small number of repositories [1].
+We choose Python for two key reasons: (1) the most popular software engineering benchmarks (e.g., SWE-bench [1]) focus on Python repositories, and (2) Python is one of the most widely used programming languages today [12].
+We then evaluated Repo2Run on this benchmark, finding it successfully built the environments for 361 repositories, achieving an 86.0% success rate, which is 63.9% higher than the leading baseline.
+Notably, Repo2Run outperforms the SWE-agent by 77.0%, demonstrating the effectiveness of the proposed agent.
+We believe Repo2Run will serve as the foundational infrastructure, driving future research at the intersection of AI and Software Engineering and facilitating the community to efficiently scale up their executable code data and environments.
+2 Formulation
+In this part, we define the task of executable environment building.
+Given a repository $R$, executable environment building aims to determine a suitable base image $B$ and a building process $\mathcal{P}$ such that the resulting environment state $S$ satisfies the verification $\varepsilon$, as defined in Equation (1):
+$\varepsilon(R,S)=0,\text{where}~{}S=\delta(B,\mathcal{P})$ (1)
+Here, $\delta$ is the state transition function, and $\varepsilon(R,S)=0$ indicates that the verification $\varepsilon$ is satisfied.
 2.1 State transition
-Environment State ($S$): The environmental state ($S$) represents the current state of the computer system, which encompasses all variables, files, cache, etc.
-Any state $S\in\mathcal{S}$, where ($\mathcal{S}$) denotes the set of all possible environment states.
-Command ($C$): The command $C$ represents an individual instruction or action that can be executed in the environment through interfaces such as bash, thus directly changing the system state.
-State Transition Function ($\delta$): The state transition function ($\delta$) defines the process through which the system transitions from one state to another state upon execution of a given command, i.e.,
-$\delta:\mathcal{S}\times\mathcal{C}\rightarrow\mathcal{S},\quad\delta(S,C)=S^{\prime}$ (1)
-Here, $S^{\prime}$ is the new state after execution of command $C$.
+$\bullet$ Environment state ($S$): The environment state $S\in\mathcal{S}$ represents the current state of the computer system, which encompasses all variables, files, cache, etc.
+$\bullet$ Command sequence ($\mathcal{C}$): The command sequence $\mathcal{C}\in\mathbb{C}$ represents a set of individual commands.
+Each individual command $C\in\mathcal{C}$ refers to an instruction or action that can be executed in the environment via interfaces like bash, thereby changing the system state.
+$\bullet$ State transition ($\delta$): The state transition $\delta$ is a function that defines the process through which the system transitions from the start state $S_{start}\in\mathcal{S}$ to the end state $S_{end}\in\mathcal{S}$ after executing of a command sequence $\mathcal{C}$, as defined in Equation (2):
+$\delta:\mathcal{S}\times\mathbb{C}\rightarrow\mathcal{S},\quad\delta(S_{start},\mathcal{C})=S_{end}$ (2)
 2.2 Base image
-Empty initial state ($S_{\emptyset}$): The empty initial state ($S_{\emptyset}$) represents a completely unconfigured, bare operating system or a purely hypothetical state without applications and configurations.
-This state is used for theoretical purposes to define the starting point of a system’s configuration.
-$S_{\emptyset}\in\mathcal{S}$.
-Base image ($B$): From the empty initial state ($S_{\emptyset}$), we can construct the base image $B$ by applying a series of actions.
-The base image $B$ is essentially the final state achieved by applying a sequence of instructions starting from $S_{\emptyset}$:
-$B=\delta^{m}(S_{\emptyset},{C^{B}_{1},C^{B}_{2},\ldots,C^{B}_{m}})$ (2)
-Here, ${C^{B}_{1},C^{B}_{2},\ldots,C^{B}_{m}}$ are the individual commands that constitute the base image.
-2.3 Configuration process and target state
-Configuration Process ($\Gamma$): The configuration process, denoted as $\Gamma$, involves transforming the environment starting from the base image $B$ by applying a series of actions.
-The result of this process is a state $S_{f}$, which we refer to as the resultant state.
-Formally, the configuration process can be represented as:
-$S_{f}=\delta^{n}\left(B,{C_{1},C_{2},\ldots,C_{n}}\right)$ (3)
-Following this transformation, it is crucial to verify whether $S_{f}$ meets all the required conditions to be considered the target state.
-This verification is performed using the state verification function.
-State verification ($\epsilon$): The state verification function ($\epsilon$) determines whether the resultant state ($S_{f}$) successfully runs all tests within the repository after executing the configuration process ($\Gamma$).
-It maps a given state ($S$) to a binary value indicating whether all tests could be run in the resultant configuration: $\epsilon(S)$ returns 0 if all tests are successfully run, and 1 otherwise.
-If $\epsilon(S_{f})=0$, it indicates that the configuration process has finished and the resultant state $S_{f}$ is considered the target state.
-Otherwise, it means that $S_{f}$ does not meet the necessary conditions.
+$\bullet$ Empty state ($S_{\emptyset}$): The empty state $S_{\emptyset}\in\mathcal{S}$ represents a completely bare operating system or a purely hypothetical state without any builds.
+$\bullet$ Base image ($B$): The base image $B\in\mathcal{S}$ is a special type of environment state, typically managed by professional teams for user convenience, such as “python:3.10”.
+Starting from the empty state $S_{\emptyset}$, the base image $B$ can be created by executing a predefined command sequence $\mathcal{C}_{B}\in\mathbb{C}$, i.e., $B=\delta(S_{\emptyset},\mathcal{C}_{B})$.
+Users can utilize these base images by simply adding their names to the Dockerfiles.
+2.3 Building process and state verification
+$\bullet$ Building process ($\mathcal{P}$): The building process $\mathcal{P}\in\mathbb{C}$ is the command sequence designed to build the environment from the base image $B$.
+We denote the resulting state as $S_{f}\in\mathcal{S}$, where $S_{f}=\delta(B,\mathcal{P})$.
+$\bullet$ State verification ($\varepsilon$): The state verification $\varepsilon$ is a Boolean function used to determine whether the resulting state $S_{f}$ successfully runs all tests in the repository $R$.
+$\varepsilon(R,S_{f})=0$ indicates that all tests in the repository $R$ can be successfully run.
+Otherwise, at least one test fails.
 3 Repo2Run
-Figure 3: An example process of Repo2Run, which illustrates two main parts: Event Stream and Repo2Run Workflow.
-The Event Stream tracks the entire action-observation process, where green boxes represent system prompts, yellow boxes represent responses from the LLM agent including the actions, and blue boxes represent observations from the Runtime Environment.
-The shaded actions indicate configurations abandoned after base image change.
-In the Event Stream, the blue text indicates that the command starts running, the green text indicates that the command runs successfully, and the red text indicates that the command fails.
-The Repo2Run Workflow consists of the internal environment and the external environment.
-The internal environment serves as the actual configuration Docker-based sandbox, which builds an actual testing runtime environment.
-The external environment executes the action-observation process and assists in the configuration process within the internal environment.
-In this part, we introduce the design of Repo2Run and atomic configuration synthesis.
-As shown in Repo2Run Workflow in Figure 3, the dual-environment architecture consists of an external environment and an internal environment.
-The internal environment is a Docker container where the actual environment configuration is executed, while the external environment assists the internal environment in setting up the configuration.
-After successfully configuring the environment, the Dockerfile generator will create the Dockerfile.
-Figure 4: An example of the Dockerfile generator to transfer the commands into a runnable Dockerfile.
+Figure 4: The workflow of Repo2Run, involving two phases: the build phase and the record phase.
+The build phase utilizes a dual-environment architecture: the internal environment with five actions for environment building, while the external environment with three actions assists the internal environment.
+The record phase converts the validated command sequence into a runnable Dockerfile for reconstructing the executable environment.
+See Appendix A for more examples of these actions.
+In this part, we introduce the design of Repo2Run.
+Given a code repository, Repo2Run synthesizes a runnable Dockerfile, enabling the building of an executable environment.
+As shown in Figure 4, this process is divided into two phases: the build phase and the record phase.
+In build phase, Repo2Run employs a dual-environment architecture consisting of an internal environment and an external environment.
+The internal environment is a Docker container where the agent can execute actions restricted to a container, e.g., test running.
+The external environment is a terminal where the agent can execute actions prohibited within a container, e.g., base image change.
+Once all tests in the internal environment are successfully executed, Repo2Run transitions to the record phase.
+In record phase, Repo2Run synthesizes a runnable Dockerfile based on the command sequence executed during the build phase.
+This Dockerfile serves as a precise record of the build process, and its execution can be viewed as a replay of the original build phase.
+Next, we introduce the external environment, internal environment, and Dockerfile synthesizer.
 3.1 External environment
-$\bullet$ Action-observation interaction: In each turn of the environment-agent interaction, the event history $E_{t}={(a_{1},o_{1}),(a_{2},o_{2}),\ldots,(a_{t-1},o_{t-1})}$ records the historical actions and observations up to time step $t$.
-Additionally, the sequence of successfully executed commands ${C_{1},C_{2},\ldots,C_{n}}$ helps remind the LLM agent of the current system state.
-At time step $t$, given the event history $E_{t}$ and the sequence of successfully executed commands $C_{1},C_{2},\dots,C_{n}$, the LLM agent generates a response including an action $a_{t}$:
-$a_{t}=\pi(E_{t},{C_{1},C_{2},\ldots,C_{n}})$ (4)
-where $\pi$ is the function of the LLM agent’s decision. The environment state then transitions based on the action $a_{t}$:
-$S_{t+1}=\delta(S_{t},a_{t})$ (5)
-This iterative process ensures that the LLM agent remains informed of the system’s current state and makes decisions accordingly, facilitating effective multi-turn interaction within the environment.
-$\bullet$ Rollback ($\rho$): As illustrated in Figure 1, commands $C_{fail}$ may lead to an uncertain state ($S_{u}$) if not successfully executed.
-When a command executes failed (i.e., the command returns a non-zero return code), the system transitions into an uncertain state as defined by:
-$S_{u}=\delta(S_{k-1},C_{fail})$ (6)
-Here, $C_{fail}$ is the failed executed command.
-If a command fails, the system will transition to the uncertain state $S_{u}$.
-To avoid such uncertainty, we introduce rollback ($\rho$).
-In the event of a failure at any step, the rollback mechanism restores the environment to the last certain state:
-$\rho:\mathcal{S}\times\mathcal{C}\rightarrow\mathcal{S},\quad\rho(S_{u},{C_{1},C_{2},\ldots,C_{k}})=S_{k-1}$ (7)
-Specifically, each time an instruction ($C$) is executed, we utilize the “docker commit” command to take a snapshot of the current state.
-Once the instruction is completed, we check its return code.
-If the return code is not 0, it indicates a failure in executing the instruction, and we replace the image with the most recently committed one.
-Besides, some instructions, such as “cat”, generally do not change the state of the environment, and thus, we do not apply the rollback mechanism to them.
-The complete list of these actions is shown in Appendix B.
-$\bullet$ Base image change: If the LLM agent identifies that the currently selected base image is incorrect or unsuitable during the environment configuration process, it can reselect the base image.
-Once a new image is chosen, all previous configurations become invalid, and the executed commands are cleared, necessitating a restart of the configuration process.
-For instance, as shown in Figure 3, when the LLM agent discovers that the “StrEnum” class cannot be imported through testing, which is only available in the Python standard library starting from version 3.11, it decides to reselect the base image to python:3.11.
-$\bullet$ Result processor: In each turn of the interaction, executing commands may generate extensive output (e.g., logs), which can greatly increase the burden of the LLM agent.
-Therefore, whenever such information is added to the event history, we truncate excessively long text, retaining only the initial and final sections up to a specific length.
-$\bullet$ Dockerfile generator: If the LLM agent successfully runs all tests in the repository in the internal environment, the configuration process needs to be recorded in a Dockerfile.
-Figure 4 illustrates the Dockerfile generator’s operation.
-4 (a) lists the internal commands actually executed in the internal environment, their return codes (e.g., 0 indicates success, otherwise failure), and the current directory (dir).
-4 (b) shows the process of generating the Dockerfile.
-The principles for the conversion process include the following steps, scanning all internal commands sequentially, and the line numbers mentioned below match the line number of Figure 4 (a):
-  1. 1.
-    By default, use python:3.10 as the base image.
-  2. 2.
-    For commands that run successfully (i.e., have a return code of 0), prepend the command with “RUN” to form a Dockerfile statement (e.g., line 3).
-Commands that fail (i.e., have a return code other than 0, such as line 4) are rolled back and not included in the Dockerfile.
-An exception is some commands typically do not change the current state (such as “cat” on line 2).
-These commands are not added to the Dockerfile.
-We show the details in Appendix B.
-  3. 3.
-    If dir is not the root directory, use cd to change to that directory before running the command, as each Dockerfile statement runs in its own directory session.
-  4. 4.
-    If a base image change statement is encountered (e.g., line 5), discard all previous configurations and switch to the new base image.
-  5. 5.
-    If a code editing command is encountered (e.g., line 9), copy the patch and editing script into the Docker container before executing the editing command.
-  6. 6.
-    If an export statement for adding environment variables is encountered (e.g., line 13), convert it to a persistent ENV statement.
-  7. 7.
-    After all statements have been scanned, check for dependency installation commands and replace any unspecified versions (e.g., “B>=1.0,<2.0” in line 11) with the actual versions downloaded in the Docker container (e.g., “B==1.5.1” in the final Dockerfile of Figure 4 (b)).
+The external environment serves as a bridge between the LLM and internal environment.
+It transmits actions from the LLM to the internal environment and relaying observations back.
+Following with the ReAct framework [13], the external environment maintains a history of thoughts, actions (i.e., successfully executed commands), and observations.
+This iterative process ensures that the LLM stays informed about the system’s state and makes accurate decisions while effectively interacting with the internal environment.
+$\bullet$ Rollback: As shown in Figure 3, commands like “commandA”  [^2] We collected some commands that can “pollute” the environment, see Appendix C.
+may “pollute” the environment, making it difficult to reproduce using a Dockerfile.
+When a command fails (i.e., returns a non-zero return code), the environment transitions to an uncertain state.
+Therefore, we introduce a rollback action to ensure the building process remains consistent.
+Specifically, before executing a command, we use “docker commit” to create a reserve snapshot of the current state (i.e., save the image at that point).
+If the return code is not 0, indicating a failure, the environment is rolled back to the most recent reserve snapshot.
+However, certain commands, referred to as safe commands (e.g., “cat”), generally do not alter the state of the environment and are exempt from rollback.
+For a complete list of safe commands, see Appendix B.2.
+$\bullet$ Base image change: If the LLM agent determines that the current base image is unsuitable during the environment building process, it can reselect and switch to a new base image.
+This change invalidates the previous building process, clearing all executed commands and requiring a restart.
+$\bullet$ Result processor: Command execution during interactions can produce extensive output (e.g., error logs), potentially overwhelming the LLM agent.
+To mitigate this, long outputs are truncated by this action, retaining only the initial and final sections up to a specified length.
 3.2 Internal environment
-As shown in the Repo2Run workflow of Figure 3, the internal environment consists of a Docker container with a base image $B$ and the code repository to be configured.
-According to the latest data [^1] https://w3techs.com/technologies/history_details/pl-python/3 from 2025, we select Python 3.10 as our default Docker base image due to its increasing usage and wide application among all Python 3 versions.
-Additionally, we categorize the executable actions within the internal environment into five types: environment monitoring, dependency installation, bash commands, code editing, and test running.
-$\bullet$ Environment monitoring: It serves as the eyes of the LLM agent in the internal environment, allowing it to observe the current state of the environment.
-These commands usually do not change the state of the environment.
-For instance, commands like “ls” and “cat” are employed to inspect directories and files.
-Additionally, commands like “pip list” and “pipdeptree [^2] https://github.com/tox-dev/pipdeptree are utilized to get the installed version of third-party libraries in the current environment.
-For example, in the fourth part of Event Stream in Figure 3, the LLM agent calls the “find” command to search for requirements.txt within the environment to determine which dependencies need to be installed.
-$\bullet$ Dependency installation: It is primarily used to install the third-party libraries necessary for running tests, including libraries managed by both “pip” and “apt-get”.
-To avoid dependency conflicts that may arise when downloading multiple packages, we implement a dependency management system in the external environment.
-Initially, the third-party libraries to be downloaded are added to a waiting list.
-If there are different constraints for the same library (e.g., “A>=>=> =1.0” and “A$<$1.0”), they will be moved to a conflict list, and the LLM agent will determine the appropriate version to download.
-We design a set of tools for the LLM agent to manage the installations.
-For instance, as shown in Figure 3, the action “waitinglist addfile” adds all elements from requirements.txt to the waiting list, while executing the “download” command, third-party libraries are taken one by one from the waiting list and installed in the internal environment.
-The installation commands (e.g., “pip install” and “apt-get install”) are executed.
-If the installation succeeds, the state of the environment is updated; if it fails, a rollback is performed.
-We show all our designed tools in Appendix C.
-$\bullet$ Test running: It detects whether the current environment is successfully configured by running “pytest” unit tests.
-If the state ($S_{f}$) of the current environment passes the tests, it indicates that the LLM agent has successfully configured the environment within the Docker container, and the internal environment’s process concludes.
-Conversely, if some tests fail, the error information is communicated back to the LLM agent, allowing it to perform further configurations and adjustments based on the errors.
-$\bullet$ Code editing: It allows the LLM agent to modify the code within the internal environment, including codes inside and outside the repository.
-Besides, to prevent the LLM agent from spuriously running the tests by directly modifying or deleting the test files, we do not permit the LLM agent to modify or delete the original test files within the repository.
-For instance, Figure 3 demonstrates code editing to correct a syntax error resulting from the improper use of double quotes in a Python f-string.
-$\bullet$ Bash commands: Aside from the commands above, we also enable the LLM agent to directly invoke bash commands to interact with the internal environment.
-As shown in Figure 3, the LLM agent uses the command “export PYTHONPATH=/repo/src”.
+The internal environment is a Docker container. Based on the latest data [14] from 2025, we select Python 3.10 as the default Docker base image due to its broadest adoption among Python versions.
+$\bullet$ Environment monitoring: It serves as the eyes of the LLM agent in the internal environment, allowing it to observe the current environment state.
+These commands typically do not change the state of the environment.
+Basic commands like “ls” and “cat” are used to inspect directories and files, “find” is used to locate files, while more advanced commands like “pip list” and “pipdeptree” [15] retrieve the versions of installed libraries and dependency relationships.
+$\bullet$ Dependency installation: It serves as installing third-party packages required for running tests, using tools like pip for Python and apt-get for system-level packages.
+To resolve potential conflicts (e.g., version constraints), a dependency management is implemented to assist the installation, where packages to be installed are added to a waiting list and resolved as needed.
+If installation succeeds, the environment is updated; otherwise, the rollback is performed.
+For detailed designs, see Appendix F.
+$\bullet$ Test running: It serves as both a compass and a checkpoint, guiding the building process and verifying whether the environment runs all tests successfully.
+By executing “pytest” for unit tests, it determines whether the Docker container has been correctly built.
+If all tests run successfully, the process concludes; otherwise, error logs are sent to the LLM agent for further adjustments.
+$\bullet$ Code editing: It enables the LLM agent to modify the code within the internal environment, including both inside and outside the repository.
+Direct code editing is rare, but sometimes needed for issues like syntax errors.
+To prevent the LLM agent from bypassing tests by directly altering or deleting test files, it is restricted from modifying or deleting the original test files within the repository.
+$\bullet$ Bash commands: Like many other LLM-based agents [2, 3, 4], Repo2Run is allowed to invoke bash commands, allowing it fully operate within the Docker container.
+It enables the agent to execute dynamically defined actions that are not pre-specified, ensuring flexibility in handling various errors.
+Figure 5: Rules for Dockerfile synthesis, illustrating how the Dockerfile synthesizer maps executed commands into Dockerfile statements using four keywords: “FROM”, “ENV”, “COPY”, and “RUN”.
+Black arrows represent the creation of statements, while green arrows indicate their transformations.
+Red text next to the arrows specifies the commands executed during each step.
+3.3 Dockerfile synthesizer
+Once the LLM agent successfully runs all tests, the Dockerfile synthesizer converts the command sequence from the building process into a runnable Dockerfile.
+It processes each command sequentially, following the rules in Figure 5 to synthesize Dockerfile statements.
+The synthesis uses four Dockerfile keywords: “FROM”, “ENV”, “COPY”, and “RUN”.
+For more details and examples, see Appendix B.
+$\bullet$ FROM: The FROM statement defines the base image and is typically the first line in a Dockerfile.
+If the LLM agent changed the base image, the FROM statement must be updated, and all subsequents are cleared to ensure consistency.
+$\bullet$ ENV: The ENV statement persistently sets environment variables in the Docker container.
+When a command with “export” is detected, it is converted into an ENV statement.
+If the vaule of an environment variable is overwritten, this statement is updated accordingly.
+$\bullet$ COPY: The COPY statement copies local files or directories into the Docker container.
+As shown in Figure 5, it is often used for code editing to import editing scripts (i.e., code_edit.py) and patches.
+$\bullet$ RUN: The RUN statement executes commands in the container, with each line creating a new bash session.
+It cannot be used for persistent environment variables.
+Specially, for every installed packages, versions are recorded to ensure reproducibility, and the RUN statement is updated accordingly.
 4 Experiment
-We evaluate the effectiveness of Repo2Run on 420 Python code repositories. As the popular option, we select gpt-4o-2024-05-13 for subsequent experiments, with the temperature uniformly set to 0.2.
+We evaluate the effectiveness of Repo2Run on 420 Python code repositories. As the popular option, we select gpt-4o-2024-05-13 for all experiments, with the temperature uniformly set to 0.2.
 4.1 Benchmark
-To the best of our knowledge, there is no prior work similar to Repo2Run that constructs an environment for arbitrary Python repositories.
-To validate the capability of Repo2Run in environment configuration, we create our new benchmark consisting of selected Python repositories from GitHub based on the following criteria:
-$\bullet$ Creation date: To minimize the impact of data contamination, we carefully selected repositories created in 2024, ensuring they are not part of the training data for mainstream large language models.
-$\bullet$ Star count: To ensure the quality of the repositories, we only select those with a star count greater than 100.
-$\bullet$ Test directory: Pytest is a leading Python testing framework that is compatible with many other Python testing tools, such as unittest.
-It identifies and runs test files and test functions that are named with a “test_” prefix or a “_test” suffix.
-Python repositories usually place their test files in the “test” or “tests” directory.
-To effectively filter repositories with unit tests, we keep repositories with the “test” or “tests” directory in their root directory to filter the repositories that most probably have unit tests.
-Through the selection process, we crawl all 449 repositories that meet all the above requirements [^3] The crawling was performed in December 2024..
-Then, we filter 420 repositories containing at least one unit test.
-These repositories constitute our benchmark for subsequent experiments.
-The details of the benchmark are presented in Appendix D.
+To the best of our knowledge, there is no prior work similar to Repo2Run that automates executable environment building.
+Existing manually constructed datasets are limited to very few repositories [1].
+To validate the capability of Repo2Run, we create a new benchmark consisting of filtered Python repositories from GitHub based on the following criteria:
+$\bullet$ Creation date: To avoid the data leakage, we select repositories created in 2024, ensuring they are not part of mainstream LLM training data.
+$\bullet$ Star count: To maintain quality, we only include repositories with more than 100 stars.
+$\bullet$ Test directory: We focus on repositories likely to contain unit tests, identified using pytest, a leading Python testing framework compatible with tools like unittest.
+Pytest detects test files with a “test_” prefix or “_test” suffix, typically located in “test” or “tests” directories.
+We only retain repositories that have these directories.
+Using these criteria, we crawled 449 repositories in December 2024 and filtered 420 containing at least one unit test to form our benchmark. For statistic of their scale, see Appendix D.
 4.2 Evaluation metrics
@@
 If the Dockerfile for a code repository successfully builds, it is regarded as a successful generation.
-Generating runnable Dockerfile is fundamental for successfully configuring the environment.
-$\bullet$ Environment Configuration Success Rate (ECSR): It represents the percentage of attempts where the method successfully configures environments.
-For a successful configuration, the generated Dockerfile must not only build successfully but also allow tests to run by “pytest” in the Docker container.
+Generating runnable Dockerfile is fundamental for successfully building the executable environment.
+$\bullet$ Environment Building Success Rate (EBSR): It represents the percentage of attempts where the method successfully builds executable environments.
+For a successful building, the generated Dockerfile must not only build successfully but also allow tests to run by “pytest” in the Docker container.
 We are only concerned with whether tests can be executed, regardless of whether they pass or fail, as outcomes of tests may inherently vary within the repository.
-Table 1: The results of different baselines.
-Metric DGSR ECSR
-pipreqs 29.8% (125) 6.0% (25)
-LLM generator 47.6% (200) 22.1% (93)
-SWE-agent 26.9% (113) 9.0% (38)
-Repo2Run 100% (420) 86.0% (361)
 4.3 Baselines
-$\bullet$ pipreqs [^4] https://github.com/bndr/pipreqs: It is an automated tool that generates a “requirements.txt” file by analyzing the import statements in the Python scripts and identifying the necessary dependencies without LLM.
+$\bullet$ pipreqs [16]: It is an automated tool that generates a “requirements.txt” file by analyzing the import statements in the Python scripts and identifying the necessary dependencies without LLM.
 Using the requirements.txt file generated by pipreqs, we create a Dockerfile.
-The detail is provided in Appendix E.
-$\bullet$ LLM generator: The “README” file in a code repository usually contains environment configuration instructions.
+The detail is provided in Appendix H.1.
+$\bullet$ LLM generator: The “README” file in a code repository usually contains environment building instructions.
 Therefore, we directly drive the LLM to read the “README” file and generate an executable Dockerfile accordingly.
-$\bullet$ SWE-agent [26]: SWE-agent establishes a custom agent-computer interface (ACI) that facilitates the LLM agent’s interaction with the repository environment by allowing actions such as reading files, editing files, and executing bash commands.
-Initially intended as an LLM agent for bug fixing, we preserve its framework and default settings, adjust its prompts, and use it as a baseline.
+$\bullet$ SWE-agent [3]: SWE-agent establishes a custom agent-computer interface (ACI) that uses the LLM agent’s interaction with the repository environment by allowing actions such as reading files, editing files, and executing bash commands.
+Initially intended as an LLM agent for bug fixing, we preserve its framework and default settings, adjust its prompts, as shown in Appendix H.2.
 4.4 Experimental Results
+Table 1: The results of different baselines, including the percentage and number of successfully generated Dockerfiles and successfully built environments.
+Metric DGSR # Successfully Generated Dockerfiles ECSR # Successfully Built Environments
+pipreqs 29.8% 125 6.0% 25
+LLM generator 47.6% 200 22.1% 93
+SWE-agent 26.9% 113 9.0% 38
+Repo2Run 100% 420 86.0% 361
+Table 2: Repository distribution across different domains and their EBSR, where SE means Software Engineering, NLP means Natural Language Processing, and DV means Data Visualization.
+Type AI / ML SE NLP DV Security Others All
+# Success 221 70 41 10 7 12 361
+# Total 267 73 47 11 8 14 420
+EBSR 82.8% 95.9% 87.2% 90.9% 87.5% 85.7% 86.0%
 The results of different baselines are presented in Table 1.
-We observe that Repo2Run consistently outperforms other baselines on both DGSR and ECSR.
-Repo2Run ultimately completed environment configuration for 361 code repositories, achieving an ECSR of 86.0%.
+We observe that Repo2Run consistently outperforms other baselines on both DGSR and EBSR.
+Repo2Run ultimately completed environment building for 361 code repositories, achieving an EBSR of 86.0%.
 It is 63.9% higher than the highest rate achieved by other methods, demonstrating great advantages.
-Due to the design of atomic configuration synthesis, Repo2Run successfully generates Dockerfiles that can be built successfully for all 420 code repositories, which other tools cannot guarantee.
+Due to the design of rollback, Repo2Run successfully generates Dockerfiles that can be built successfully for all 420 code repositories, which other actions cannot guarantee.
+For details of each repostories, see Appendix G.
 For pipreqs, the main failures come from two reasons.
@@
 This happens in 30 repositories (7.1%).
-Second, even when requirements.txt is generated, it might not download properly due to package version conflicts.
+Second, even when requirements.txt is generated, it might not be downloaded properly due to package version conflicts.
 This occurs in 265 repositories (63.1%).
@@
 This indicates that a general agent framework cannot guarantee the generation of runnable Dockerfiles.
-Ensuring mechanisms like atomic configuration synthesis are necessary to effectively use the interactive information from the agent to generate runnable Dockerfiles.
-Table 2: The results of ablation experiments.
-Metric DGSR ECSR
-w/o dual-environment 92.4% (388) 41.7% (175)
-w/o Dockerfile generator 19.5% (82) 13.8% (58)
-Repo2Run 100% (420) 86.0% (361)
+Ensuring mechanisms like rollback is necessary to effectively use the interactive information from the agent to generate runnable Dockerfiles.
+As shown in Table 2, we manually categorized the 420 repositories into six domains.
+EBSR across these domains all exceed 80%, outperforming all baselines and demonstrating Repo2Run’s consistent performance across different domains.
+Prior research [17, 18] shows that the real-world benchmarks differs from synthetic ones and better reflect a model’s true capabilities.
+Our benchmark is built entirely from real-world GitHub data, filtered only by repository creation date, star count, and test, with no additional filtering.
+Thus, it accurately represents the current distribution of Python repositories.
+Moreover, a previous real-world benchmark [18] includes 59.6% of AI/ML repositories, closely matching the 63.6% proportion in our benchmark.
+This alignment further confirms that our benchmark reflects the real-world scenarios.
 4.5 Ablabtion of Repo2Run
-To investigate the impacts of the dual-environment architecture and Dockerfile generator separately, as two parts of atomic configuration synthesis, we separately remove each component of them.
-For the experiment without the dual-environment architecture, we retain only the internal environment’s bash commands as the most basic interface and remove all other tools.
-For the experiment without the Dockerfile generator, we directly instruct the LLM to generate a runnable Dockerfile.
+Table 3: The results of ablation experiments, including the percentage and number of successfully generated Dockerfiles and successfully built environments.
+Metric DGSR # Successfully Generated Dockerfiles ECSR # Successfully Built Environments
+w/o dual-environment 92.4%  $\downarrow\,7.6$ 388 41.7%  $\downarrow\,44.3$ 175
+$\hookrightarrow$ w/o rollback 96.9%  $\downarrow\,3.1$ 407 83.6%  $\downarrow\,2.4$ 351
+w/o Dockerfile generator 19.5%  $\downarrow\,80.5$ 82 13.8%  $\downarrow\,72.2$ 58
+Repo2Run 100% 420 86.0% 361
+To investigate the impacts of the dual-environment architecture and Dockerfile synthesizer separately, as two parts of Repo2Run, we separately remove each component of them.
+For the experiment without the dual-environment architecture, we retain only the internal environment’s bash commands as the most basic interface and remove all other actions.
+For the experiment without the Dockerfile synthesizer, we directly instruct the LLM to synthesize a runnable Dockerfile.
 Experimental result of the ablation study is shown in Table 3.
@@
 The main reason for this drop is the removal of rollback and other designs, making the system more prone to entering uncertain states and subsequently failing to reproduce.
-In addition, ECSR shows a 44.3% decrease, primarily because the simplification of design makes it more difficult for the LLM agent to complete configuration in the internal environment.
-Besides, removing the Dockerfile generator directly leads to an 80.5% drop in DGSR.
+In addition, EBSR shows a 44.3% decrease, primarily because the simplification of design makes it more difficult for the LLM agent to execute all tests in the internal environment.
+Besides, removing the Dockerfile synthesizer directly leads to an 80.5% drop in DGSR.
 This indicates that having the LLM directly generate Dockerfiles is unlikely to fully follow the event history, resulting in Dockerfiles that fail to build successfully.
-This also directly causes a sharp decline in ECSR.
-It is also observed that Repo2Run without the Dockerfile generator is outperformed by the LLM generator.
-This is because the LLM generator leverages the “README” file, which provides a clear, simple and high-level overview of the environment configuration, allowing for more accurate Dockerfile generation.
-In contrast, the event history-based approach lacks this context, making it harder for the LLM to fully understand the configuration goals.
-However, the Dockerfile generator effectively utilizes the detailed event history, highlighting the complementary roles of both components of Repo2Run in generating a reliable Dockerfile.
+This also directly causes a sharp decline in EBSR.
+Specially, we also perform an ablation study for the rollback.
+The results show that removing the rollback causes 3.1% of generated Dockerfiles to become unrunnable, demonstrating the effectiveness of the rollback mechanism.
+It is also observed that Repo2Run without the Dockerfile synthesizer is outperformed by the LLM generator.
+This is because the LLM generator leverages the “README” file, which provides a clear, simple and high-level overview of the executable environment building, allowing for more accurate Dockerfile synthesizer.
+In contrast, the event history-based approach lacks this context, making it harder for the LLM to fully understand the building goals.
+However, the Dockerfile synthesizer effectively utilizes the detailed event history, highlighting the complementary roles of both components of Repo2Run in generating a reliable runnable Dockerfile.
 5 Discussion
-5.1 Tool usage frequency
-Figure 5: Tool usage frequency of the success configuration.
-As shown in Figure 5, we analyze the invocation times of various action types in 361 successfully configured projects, including the five actions within the internal environment and the action of changing the base image.
-Bash commands are the most frequently used action, as they encompass the majority of instructions.
-Additionally, we observe that the LLM agent tends to call dependency installation quite frequently, averaging about 18 times per configuration, which means roughly 18 dependencies are installed per configuration on average.
-Moreover, the LLM agent calls test running approximately 3.5 times per configuration on average, which typically helps the agent better identify issues.
-Among successful configurations, we see 48 instances of changing the base image, accounting for 13.3%  of success cases, indicating that the initial selection of the base image is often incorrect and requires subsequent adjustments.
+Figure 6: Time distribution of successful building.
 5.1 Time consumption
-Figure 6: The time distribution of each successful configuration.
-Figure 6 shows the distribution of time spent successfully configuring 361 code repositories.
-The average time for successful configuration using Repo2Run is 29.03 minutes.
-111 (30.7%) of the repositories are successfully configured in less than 10 minutes.
-Additionally, our empirical study through sampling indicates an average manual configuration time of 21.33 minutes (Appendix A).
-Considering network differences and randomness, Repo2Run achieves a time consumption comparable to manual configuration.
-Additionally, for complex issues, Repo2Run shows greater advantages over manual configuration.
-Repo2Run successfully configures all the cases that were manually successful in our empirical study.
-Moreover, the cases where Repo2Run fails are also not successfully configured manually.
-5.3 Case study
-For the code repositories that fail to configure, we manually inspect the reasons for failures and find that most are due to issues within the repositories themselves.
-Figure 7 illustrates a “module not found” error when configuring the repository “jialuechen/deepfolio”.
-In this case, the issue arises from the absence of updating the unit tests.
-The test file “test_stats.py” attempts to import modules from “deepfolio.stats”, but “stats” does not exist in “deepfolio”.
-Consequently, no matter how the LLM agent operates, it cannot directly run this test.
-This highlights the importance for developers to continuously update existing unit tests as the code repository evolves.
+Figure 6 shows the distribution of time spent successfully building 361 code repositories.
+The average time for successful building using Repo2Run is 29.03 minutes.
+111 (30.7%) of the repositories are successfully built in less than 10 minutes.
+Additionally, our empirical study through sampling indicates an average manual building time of 21.33 minutes (Appendix E).
+Considering network differences and randomness, Repo2Run achieves a time consumption comparable to manual building.
+Additionally, for complex issues, Repo2Run shows greater advantages over manual building.
+Repo2Run successfully builds all the cases that were manually successful in our empirical study.
+Moreover, the cases where Repo2Run fails are also not successfully built manually.
+Table 4: Analysis of failed cases in Repo2Run.
+Category # Case (%)
+Hardware Insufficiency 21 (35.6%)
+Missing Token 1 (1.7%)
+Repository Defects 5 (8.5%)
+Dependency Installation Timeout 8 (13.5%)
+Runtest Timeout 24 (40.7%)
+Table 5: Comparison of Dockerfile quality. “Warning” counts in each group show Repo2Run vs. baseline.
+Baseline # “Warning”
+vs. pipreqs 5.17 / 6.0
+vs. LLM generator 5.42 / 10.16
+vs. SWE-agent 5.43 / 6.49
+5.2 Failure case study
+For the code repositories that fail to build, we manually inspect the reasons for failures and find that most are due to issues within the repositories themselves.
+Table 4 demonstrates the analysis of 59 failed cases in Repo2Run.
+The first three categories (45.8%) primarily fail due to external factors, while the latter two categories (54.2%) are unable to be configured successfully within the allotted time due to excessive dependencies and extensive tests inherent to the repositories.
+We also provide a detailed case in Appendix I.
+5.3 Dockerfile Quality
+We utilize hadolint [19] to analyze the quality of the Dockerfile generated by Repo2Run, comparing with Dockerfiles successfully generated by three baselines and counting the number of warnings.
+Fewer warnings indicate higher quality in Dockerfile generation.
+As shown in Table 5, Dockerfiles generated by Repo2Run’s consistently have fewer warnings than those generated by baselines.
+Specifically, compared to pipreqs, LLM generator, and SWE-agent, Repo2Run reduces warnings from 6.0 to 5.17, 10.16 to 5.42, and 6.49 to 5.43, respectively.
+The results indicate that Repo2Run  produces higher-quality Dockerfiles with fewer potential issues.
+5.4 Limitation
+Repo2Run can automate building executable environments for many repositories but may still encounter challenges in some edge cases.
+For complex environments that remain unresolved, manual verification and intervention are necessary.
+Nonetheless, as an important step toward full automation, we believe this work makes a significant contribution to the AI community’s infrastructure.
 6 Related Work
-6.1 Environment configuration and Dockerfile generation
+6.1 Environment building and Dockerfile generation
 Oss-Fuzz-Gen [20] relies on predefined build instructions (such as “./bootstrap.sh”, “./configure”, “make”) to build projects for fuzzing but lacks of flexibility for diverse projects when specified files are absent.
-Existing solutions that help developers write Dockerfiles broadly fall into three categories: (1) Template-based generators that create Dockerfiles based on project context [2, 18], (2) Task-specific tools such as DockerizeMe which supports environment dependency inference for Python projects  [11] and DockerGen  [27] for dependency recommendations based on knowledge graphs built from existing Dockerfiles, and (3) Code completion tools including GitHub Copilot [7] and HumpBack [8] that generate suggestions for developers while writing Dockerfiles.
+Existing solutions that help developers write Dockerfiles broadly fall into three categories: (1) Template-based generators that create Dockerfiles based on project context [21, 22], (2) Task-specific tools such as pipreqs [16], DockerizeMe [23] which supports environment dependency inference for Python projects and DockerGen  [24] for dependency recommendations based on knowledge graphs built from existing Dockerfiles, and (3) Code completion tools including GitHub Copilot [5] and HumpBack [25] that generate suggestions for developers while writing Dockerfiles.
 The use of deep learning models to generate Dockerfiles based on natural language specifications of software requirements is also investigated [26].
@@
 To extend capabilities beyond interactive dialogue, the action component employs various external tools [41, 42], such as searching tools, file operations, and GUI operations.
-Nowadays, LLM-based agents have demonstrated superior performance compared to standalone LLMs in various software engineering tasks [13, 5, 1].
-However, no LLM-based agents are specifically designed for environment configuration currently.
-To fill this gap, this paper employs a novel approach for automated coding environment configuration and Dockerfile generation.
-Figure 7: An example of “module not found” error due to the absence of updating the unit tests in the repository.
+Nowadays, LLM-based agents have demonstrated superior performance compared to standalone LLMs in various software engineering tasks [43, 44, 45, 46, 47].
+However, no LLM-based agents are specifically designed for executable environment building currently.
+To fill this gap, this paper employs a novel approach for automated executable coding environment building and runnable Dockerfile generation.
 7 Conclusion
-In this paper, we propose Repo2Run, the first LLM-based agent for automated coding environment configuration and Dockerfile generation for Python repositories based on atomic configuration synthesis.
-With a dual-environment architecture and a Dockerfile generator, Repo2Run is able to select and change the base image, manage and install dependencies based on action observation and rollback mechanism, and utilize bash commands and existing test suites.
+In this paper, we propose Repo2Run, the first LLM-based agent for automated coding executable environment building and Dockerfile generation for Python repositories.
+With a dual-environment architecture and a Dockerfile synthesizer, Repo2Run is able to select and change the base image, manage and install dependencies based on action observation and rollback mechanism, and utilize bash commands and existing test suites.
 Our evaluation of 420 popular Python repositories hosted on GitHub demonstrates the effectiveness of Repo2Run with an 86.0% success rate.
-Impact Statement
-Repo2Run demonstrates the potential and impact that LLM-based agents can have towards automating coding environment configuration and docker file generation.
-We are hopeful that Repo2Run may inspire more works that advance broader aspects of automated software engineering.
-There might be more potential societal consequences of our work, none of which which we feel must be specifically highlighted here.
+We believe Repo2Run will serve as the foundational infrastructure, enabling the community to efficiently scale up executable environments.
 Appendix
-Appendix A Manual experiment
-A.1 Settings
+Appendix A Repo2Run details
+A.1 Example of Repo2Run
+Figure 7 shows an example process of Repo2Run, including the event stream and Repo2Run  workflow. Below, we introduce the executed action based on the example:
+  • Base Image Change: In the tenth part of Event Stream, when the LLM agent discovers that the “StrEnum” class cannot be imported through testing, which is only available in the Python standard library starting from version 3.11, it decides to reselect the base image to python:3.11.
+  • Environment monitoring: In the fourth part, the LLM agent calls the “find” command to search for requirements.txt within the environment to determine which dependencies to install.
+  • Dependency installation: In the sixth part, the action “waitinglist addfile” adds all elements from requirements.txt to the waiting list, while executing the “download” command.
+Third-party libraries are taken one by one from the waiting list and installed in the internal environment.
+The installation commands (e.g., “pip install” and “apt-get install”) are executed.
+  • Code editing: The eighteenth part demonstrates code editing to correct a syntax error resulting from the improper use of double quotes in a Python f-string.
+  • Bash commands: The sixteenth part shows that the LLM agent uses the command “export PYTHONPATH=/repo/src”.
+  • Test running: The eighth part shows that the LLM agent calls the “runtest” command to execute unit tests.
+Figure 7: An example process of Repo2Run, which illustrates two main parts: Event Stream and Repo2Run Workflow.
+The Event Stream tracks the entire action-observation process, where the green box represents the abbreviated system prompt, yellow boxes represent responses from the LLM agent including the actions, and blue boxes represent observations from the Runtime Environment.
+The shaded actions indicate buildings abandoned after base image change.
+In the Event Stream, the blue text indicates that the command starts running, the green text indicates that the command runs successfully, and the red text indicates that the command fails.
+The Repo2Run Workflow consists of the internal environment and the external environment.
+The internal environment serves as the actual buildings Docker-based sandbox, which builds an actual testing runtime environment.
+The external environment executes the action-observation process and assists in the building process within the internal environment.
+A.2 Action usage frequency
+Figure 8: Action usage frequency of the success building.
+As shown in Figure 8, we analyze the invocation times of various action types in 361 successfully built projects, including the five actions within the internal environment and the action of base image change.
+Bash commands are the most frequently used action, as they encompass the majority of instructions.
+Additionally, we observe that the LLM agent tends to call dependency installation quite frequently, averaging about 18 times per building, which means roughly 18 dependencies are installed per building on average.
+Moreover, the LLM agent calls test running approximately 3.5 times per building on average, which typically helps the agent better identify issues.
+Among successful buildings, we find 48 instances of changing the base image, accounting for 13.3%  of success cases, indicating that the initial selection of the base image is often incorrect and requires subsequent adjustments.
+Appendix B Dockerfile synthesizer details
+B.1 Example
+Figure 9: An example of Dockerfile synthesizer to transfer the commands into a runnable Dockerfile.
+If the LLM agent successfully runs all tests in the repository in the internal environment, the building process needs to be recorded in a Dockerfile.
+Figure 9 illustrates the Dockerfile synthesizer’s operation.
+Figure 9 (a) lists the internal commands actually executed in the internal environment, their return codes (e.g., 0 indicates success, otherwise failure), and the current directory (dir).
+Figure 9 (b) shows the process of generating the Dockerfile.
+The principles for the conversion process include the following essential steps: scanning all internal commands sequentially, and the line numbers mentioned below precisely match the line number of Figure 9 (a):
+  1. 1.
+    By default, use python:3.10 as the base image.
+  2. 2.
+    For commands that run successfully (i.e., have a return code of 0), prepend the command with “RUN” to form a Dockerfile statement (e.g., line 3).
+Commands that fail (i.e., have a return code other than 0, such as line 4) are rolled back and not included in the Dockerfile.
+An exception is that some commands typically do not change the current state (such as “cat” on line 2).
+They are not added to the Dockerfile.
+We show the complete list in Appendix B.2.
+  3. 3.
+    If dir is not the root directory, use cd to change to that directory before running the command, as each Dockerfile statement runs in its own directory session.
+  4. 4.
+    If a base image change statement is encountered (e.g., line 5), discard all previous buildings and switch to the new base image.
+  5. 5.
+    If a code editing command is encountered (e.g., line 9), copy the patch and editing script into the Docker container before executing the editing command.
+  6. 6.
+    If an export statement for adding environment variables is encountered (e.g., line 13), convert it to a persistent ENV statement.
+  7. 7.
+    After all statements have been scanned, check for dependency installation commands and replace any unspecified versions (e.g., “B>=1.0,<2.0” in line 11) with the actual versions downloaded in the Docker container (e.g., “B==1.5.1” in the final Dockerfile of Figure 9 (b)).
+B.2 Safe command
+If Repo2Run executes the following commands without using “$>$” or “>>much-greater-than$>$ >” for output redirection, they are regarded to be safe commands that typically do not affect the system.
+Therefore, rollback is not necessary, and they are not added to the generated Dockerfile.
+[“cd”, “ls”, “cat”, “echo”, “pwd”, “whoami”, “who”, “date”, “cal”, “df”, “du”, “free”, “uname”, “uptime”, “w”, “ps”, “pgrep”, “top”, “dmesg”, “tail”, “head”, “grep”, “find”, “locate”, “which”, “file”, “stat”, “cmp”, “diff”, “xz”, “unxz”, “sort”, “wc”, “tr”, “cut”, “paste”, “tee”, “awk”, “env”, “printenv”, “hostname”, “ping”, “traceroute”, “ssh”]
+Appendix C Pollution commands
+Figure 10 shows two practical examples.
+As shown in Figure 10 (a), when we run “pip install cupy”, the command fails to execute because the CUDA environment is invalid.
+However, it still introduces the “fastrlock” and “numpy” packages which is not present in the original environment.
+In Figure 10 (b), when we run the “rm -rf /path/to/logs” command to delete the directory, some files fail to be deleted due to a lack of permission.
+However, the files that have already been deleted (e.g., log1.txt and log2.txt) are not restored.
+Such “pollution” caused by failed commands may make the environment enter the uncertain state.
+Figure 10: Examples of failed command causing “pollution” to the environment.
+We also identified 30 Python packages that may introduce “pollution” when installing via pip, as listed in Table LABEL:tab:pollution_example.
+This process can be easily reproduced in a clean Docker environment (such as the initial environment shown in Figure 10 (a)).
+For example, simply running “pip install adb” — even without specifying a version, which defaults to the latest — can cause such “pollution”.
+This demonstrates that such scenarios are quite common in real-world development and also underscores the importance of the rollback action in Repo2Run.
+Table 6: 30 examples of "pollution" during pip installation.
+“Package_name” represents the name of the package that failed to download, "pollution" represents other packages introduced by the failed download, and "pollution" count represents the number of these introduced packages.
+package_name "pollution" "pollution" count
+zbarlight pillow 1
+mxnet-cu91 chardet, idna, urllib3 3
+texthero blis, catalogue, certifi, charset-normalizer, click, … 41
+url six 1
+robotframework-ride certifi, charset-normalizer, idna, numpy, packaging, … 10
+adb libusb1, typing 2
+onegov-core fastcache, mailthon, passlib, polib, pytz, … 12
+postal six 1
+changes requests 1
+mxnet-cu75mkl chardet, idna, urllib3 3
+winpdb numpy, six 2
+slybot attrs, Automat, certifi, chardet, charset-normalizer, … 30
+fbprophet aiohappeyeyeballs, aiohttp, aiosignal, appdirs, async-timeout, … 44
+mxnet-cu75 chardet, idna, urllib3 3
+libarchive nose 1
+atari-py numpy, six 2
+reppy cachetools, certifi, charset-normalizer, idna, python-dateutil, … 8
+sovrin leveldb, libnacl, msgpack-python, orderedset, Pympler, … 15
+scrapely numpy, six, w3lib 3
+kevinsr python-version 1
+pysurvive colored, numpy, pillow, psutil, pygtrie, … 6
+horovod cffi, cloudpickle, packaging, psutil, PyYAML, … 6
+cupy fastrlock, numpy 2
+face-recognition face-recognition-models 1
+nsot gunicorn, idna, ipaddress, ipython, itypes, … 15
+gooey colored, numpy, pillow, psutil, pygtrie, … 6
+wxpython numpy, six 2
+neuralcoref annotated-types, blis, boto3, botocore, catalogue, … 47
+apache-airflow-backport-providers-apache-hive apispec, argcomplete, attrs, babel, cached-property, … 58
+bcolz numpy 1
+Appendix D Benchmark statistic
+As demonstrated in Table LABEL:tab:loc_staticstic, we list the statistics on lines of code for repositories in our constructed benchmark.
+Table 7: Statistics on Lines of Code (LOC) for repositories in our benchmark.
+The table below illustrates the statistics regarding the lines of code (LOC) of all repositories in the benchmark.
+It separately counts the number of repositories with LOC greater than 10,000 and greater than 100,000.
+The last two rows provide reference LOC for well-known industry-level repositories.
+Configuration state #LOC > 10,000 #LOC > 100,000 Medium Max
+Success(361) 220 (60.9%) 27 (7.5%) 14,578 6,431,084
+Fail(59) 50 (84.7%) 12 (30.6%) 43,001 3,584,542
+All(420) 270 (64.3%) 39 (9.3%) 15,854 6,431,084
+requests (reference) Yes No 12,782 12,782
+pandas (reference) Yes Yes 571,849 571,849
+Appendix E Empirical study
+E.1 Manual Experiment
+E.1.1 Settings
 To ensure a fair comparison among participants, we conducted training and demonstrated examples before the experiment to ensure everyone understood the procedure.
 Additionally, to minimize discrepancies in time consumption due to network factors, all participants conducted the experiment in the same network environment.
-A.2 Survey
+E.1.2 Survey
 We selected eight technical staff from internet companies to participate in the experiment and conducted a survey regarding their backgrounds prior to the experiment.
@@
 Seven participants have experience in complex development projects, while one has experience in multiple small-scale projects.
-Regarding environment configuration, three participants indicated that they spend a significant amount of time configuring the environment when faced with an unfamiliar code repository; another three stated that, although they spend a long time, it is generally manageable; two participants reported spending minimal time.
-In terms of successful environment configuration in their regular work, three participants mentioned they only fail with extremely complex environments, while five indicated they can build environments for most medium-scale repositories.
-As for their confidence in successfully configuring unfamiliar environments, six participants expressed that they are usually successful, and two said they are sometimes successful.
-When it comes to the amount of time they are willing to wait to configure an unfamiliar repository, four participants are willing to wait for over 90 minutes, two are willing to wait 40-60 minutes, one is willing to wait 20-40 minutes, and one is only willing to wait 10-20 minutes.
-Participants’ overall evaluation of configuring code running environments is diverse: one finds it very troublesome with many issues, four find it somewhat troublesome, two indicate moderate difficulty, and one finds it relatively simple.
-Additionally, all participants expressed a high or very high willingness to use a tool that could automatically configure the environment for an unfamiliar repository.
-A.3 Experiment guideline
+Regarding environment building, three participants indicated that they spend a significant amount of time building the environment when faced with an unfamiliar code repository; another three stated that, although they spend a long time, it is generally manageable; two participants reported spending minimal time.
+In terms of successful environment building in their regular work, three participants mentioned they only fail with extremely complex environments, while five indicated they can build environments for most medium-scale repositories.
+As for their confidence in successfully building unfamiliar environments, six participants expressed that they are usually successful, and two said they are sometimes successful.
+When it comes to the amount of time they are willing to wait to build an unfamiliar repository, four participants are willing to wait for over 90 minutes, two are willing to wait 40-60 minutes, one is willing to wait 20-40 minutes, and one is only willing to wait 10-20 minutes.
+Participants’ overall evaluation of building code running environments is diverse: one finds it very troublesome with many issues, four find it somewhat troublesome, two indicate moderate difficulty, and one finds it relatively simple.
+Additionally, all participants expressed a high or very high willingness to use a tool that could automatically build the environment for an unfamiliar repository.
+E.1.3 Experiment guideline
 We request all participants to conduct the experiment following the guidelines below:
 1. Environment Setup
-Configure the Docker environment.
+Build the Docker environment.
 Verify the installation is successful: If installed correctly, you should be able to use the following command (python:3.10 is just an example; select the base image according to your requirements):
@@
 2. Overall Procedure
-Our objective is to install the given package in a Docker container, configure its environment, and be able to run its internal tests.
-During the process, record the time developers spend configuring the environment and eventually save the logs using docker logs.
-2.1 Review the Repository to be Configured (Optional)
-Review the GitHub repository that needs to be configured.
+Our objective is to install the given package in a Docker container, build its environment, and be able to run its internal tests.
+During the process, record the time developers spend building the environment and eventually save the logs using docker logs.
+2.1 Review the Repository to be Built (Optional)
+Review the GitHub repository that needs to be built.
 2.2 Determine the Docker Base Image (Generally start with python:3.10)
@@
 2.5 Download the Repository
-Select the repository to be configured and download its GitHub repository to a location (generally directly in the root directory):
+Select the repository to be built and download its GitHub repository to a location (generally directly in the root directory):
 git clone https://github.com/{full_name}.git
-2.6 Configure the Environment
-Now, use your skills to configure: First, enter the downloaded file directory, for example:
+2.6 Build the Environment
+Now, use your skills to build: First, enter the downloaded file directory, for example:
 cd wddbfs
@@
 pytest --collect-only -q
-At this point, you can use your experience and information from the repository documentation and debugging error messages to configure. However, there are a few restrictions:
+At this point, you can use your experience and information from the repository documentation and debugging error messages to build. However, there are a few restrictions:
 Do not directly edit the test files! (Files starting with test_ or ending with _test.py). Do not directly delete test files! Editing the original repository files is not recommended.
@@
 If pytest completes successfully, the task is done.
-Scenario Two: If you feel the package is extremely difficult to configure, for example exceeding your patience threshold (refer to your usual development habits), you may also terminate.
+Scenario Two: If you feel the package is extremely difficult to build, for example exceeding your patience threshold (refer to your usual development habits), you may also terminate.
 Once finished, input exit to exit.
@@
 You need to fill out the form according to your feelings.
+Here, Table LABEL:tab:form presents the form that developers are required to complete after the environment building process, including items such as “Is it successful?”, “Final base image used”, “Reason for failure”, and so on.
+Besides, Tables LABEL:tab:form_example1 and LABEL:tab:form_example2 show two examples filled out by developers.
+Table 8: The form that developers should fill in after the environment building.
 Question Description
@@
 Does not include time spent on decision-making and research.
-Provide an approximate range: ¡3 minutes 3-5 minutes 6-10 minutes 11-20 minutes 20-40 minutes 40-60 minutes 60-90 minutes 90+ minutes
-Longest time-consuming process Describe the most time-consuming configuration process, such as: Downloading a specific dependency Resolving a specific error Incorrect Python version selection
+Provide an approximate range: <3 minutes 3-5 minutes 6-10 minutes 11-20 minutes 20-40 minutes 40-60 minutes 60-90 minutes 90+ minutes
+Longest time-consuming process Describe the most time-consuming building process, such as: Downloading a specific dependency Resolving a specific error Incorrect Python version selection
 Tolerance level Based on your subjective feeling and development experience.
-Rate the process (1-5): 1: Extremely unbearable 2: Somewhat unbearable, but manageable 3: Neutral, tolerable 4: Comfortable, no significant discomfort 5: Very comfortable, highly satisfactory configuration experience
-Configuration difficulty Based on the complexity of the configuration process (ignoring time spent): Rate the process (1-5): 1: Very simple, completed with intuition and experience, no reference materials needed 2: Fairly simple, referred to basic materials (e.g., README), simple overall steps 3: Moderate difficulty, encountered some issues, but manageable 4: Difficult, required extensive debugging and configuration 5: Very difficult, needed numerous references and encountered unresolved or time-consuming issues
-Materials referenced List the materials used for reference (e.g., README, internal configuration files, online searches, GPT). If none, mention “Directory”.
-Biggest challenge during the process Describe the most troublesome aspect of the configuration process, such as long wait times, unclear error messages, dependency version conflicts, inability to find required software versions, etc.
+Rate the process (1-5): 1: Extremely unbearable 2: Somewhat unbearable, but manageable 3: Neutral, tolerable 4: Comfortable, no significant discomfort 5: Very comfortable, highly satisfactory building experience
+Building difficulty Based on the complexity of the building process (ignoring time spent): Rate the process (1-5): 1: Very simple, completed with intuition and experience, no reference materials needed 2: Fairly simple, referred to basic materials (e.g., README), simple overall steps 3: Moderate difficulty, encountered some issues, but manageable 4: Difficult, required extensive debugging and building 5: Very difficult, needed numerous references and encountered unresolved or time-consuming issues
+Materials referenced List the materials used for reference (e.g., README, internal building files, online searches, GPT). If none, mention “Directory".
+Biggest challenge during the process Describe the most troublesome aspect of the building process, such as long wait times, unclear error messages, dependency version conflicts, inability to find required software versions, etc.
 Example 1:
+Table 9: A sample form filled by a developer after environment building, illustrating a successful environment building.
 Question Answer
@@
 Tolerance level 5
-Configuration difficulty 1
+Building difficulty 1
 Materials referenced Directory
-Biggest challenge during the process Waiting for configuration and installation
+Biggest challenge during the process Waiting for building and installation
 Example 2:
+Table 10: A sample form filled by a developer after environment building, illustrating a failed environment building.
 Question Answer
@@
 Waiting time 40-60 minutes
-Longest time-consuming process Resolving FileNotFoundError, ImportError, incorrect Python version selection
+Longest time-consuming process Resolving a series of FileNotFoundError, ImportError, incorrect Python version selection
 Tolerance level 2
-Configuration difficulty 5
+Building difficulty 5
 Materials referenced README, GPT, StackOverflow
 Biggest challenge during the process Dependency version conflicts, long wait times
-A.4 Repository assignment
-We randomly assigned each participant four unique code repositories that were successfully configured by Repo2Run.
-Additionally, each participant was assigned two code repositories that Repo2Run failed to configure.
+E.1.4 Repository assignment
+We randomly assigned each participant four unique code repositories that were successfully built by Repo2Run.
+Additionally, each participant was assigned two code repositories that Repo2Run failed to build.
 To avoid chance occurrences, each failed repository was assigned to two different participants.
 Below is the list of selections:
-Successfully configured:
+Successfully built:
 [alexwlchan/safari-webarchiver, ManiMozaffar/aioclock, mixedbread-ai/batched, mobiusml/gemlite, circlemind-ai/fast-graphrag, knowsuchagency/promptic, mbodiai/embodied-agents, modelscope/agentscope, Adibvafa/CodonTransformer, kennethreitz/simplemind, lmstudio-ai/venvstacks, mlecauchois/micrograd-cuda, IST-DASLab/PanzaMail, MetaGLM/zhipuai-sdk-python-v4, openai/mle-bench, RealOrangeOne/django-tasks, basf/MolPipeline, dai-motoki/zoltraak, lucidrains/alphafold3-pytorch, mistralai/mistral-common, BMPixel/moffee, DataformerAI/dataformer, jahwag/ClaudeSync, volfpeter/htmy, Genentech/gReLU, OpenNLPLab/lightning-attention, paradigmxyz/spice, reagento/dishka, arcee-ai/fastmlx, KyanChen/RSMamba, neuralmagic/guidellm, simonw/files-to-prompt]
-Failed to configure:
+Failed to build:
 [zhuqinfeng1999/Samba, dongxuyue/Open-ReplaceAnything, LazyAGI/LazyLLM, jialuechen/deepfolio, KOSASIH/pi-nexus-autonomous-banking-network, AARG-FAN/Yolo_for_Wukong, plinder-org/plinder, expectedparrot/edsl]
-A.5 Results
-Based on the times shown in the logs, we calculated that the average configuration time for each repository is 21.33 minutes.
-Furthermore, none of the repositories that Repo2Run failed to configure were successfully configured manually.
-Additionally, in the manual experiment, five environments that were successfully configured by Repo2Run were not successfully configured, representing 15.6% of the total successfully configured assignments.
-Appendix B Safe commands
-If Repo2Run executes the following commands without using “$>$” or “>>much-greater-than$>$ >” for output redirection, they are regarded to be safe commands that typically do not affect the system.
-Therefore, rollback is not necessary, and they are not added to the generated Dockerfile.
-[“cd”, “ls”, “cat”, “echo”, “pwd”, “whoami”, “who”, “date”, “cal”, “df”, “du”, “free”, “uname”, “uptime”, “w”, “ps”, “pgrep”, “top”, “dmesg”, “tail”, “head”, “grep”, “find”, “locate”, “which”, “file”, “stat”, “cmp”, “diff”, “xz”, “unxz”, “sort”, “wc”, “tr”, “cut”, “paste”, “tee”, “awk”, “env”, “printenv”, “hostname”, “ping”, “traceroute”, “ssh”]
-Appendix C Repo2Run tools
+E.1.5 Results
+Based on the times shown in the logs, we calculated that the average building time for each repository is 21.33 minutes.
+Furthermore, none of the repositories that Repo2Run failed to build were successfully built manually.
+Additionally, in the manual experiment, five environments that were successfully built by Repo2Run were not successfully built, representing 15.6% of the total successfully built assignments.
+The table information and logs are too lengthy, so we place them on https://anonymous.4open.science/r/Repo2Run.
+E.2 Industry-level repository building validation
+As shown in Table LABEL:tab:industry_evaluation, we present the evaluation of Repo2Run  on 59 popular industry-grade repositories.
+Additionally, Table LABEL:tab:industry_evaluation1 shows the successful configuration of additional industry-grade repositories.
+Table 11: Evaluation of Repo2Run on 59 popular industry-grade repositories. The repositories were selected from GitHub Ranking [48], filtering out those without tests and with #LOC < 10,000.
+Configuration state #LOC > 10,000 #LOC > 100,000 Medium Max
+Success(30) 30 (100%) 14 (46.7%) 96,157 2,389,065
+Fail(29) 29 (100%) 24 (82.6%) 333,464 2,669,973
+All(59) 59 (100%) 40 (9.3%) 124,352 2,669,973
+Table 12: Successful additional configuration of 30 industry-grade repositories.
+Repository #LOC Repository #LOC Repository #LOC
+comfyanonymous/ ComfyUI 737,543 OpenBB-finance/OpenBB 2,389,065 scrapy/scrapy 77,278
+yt-dlp/yt-dlp 155,662 keras-team/keras 204,537 AUTOMATIC1111/ stable-diffusion-webui 40,448
+pallets/flask 25,190 geekan/MetaGPT 82,305 psf/black 118,551
+deepfakes/faceswap 74,690 streamlit/streamlit 201,894 psf/requests 12,782
+lllyasviel/Fooocus 328,199 labmlai/annotated_ deep_learning _paper_implementations 510,851 RVC-Boss/GPT-SoVITS 34,658
+mingrammer/ diagrams 11,994 OpenInterpreter/open-interpreter 15,818 hiyouga/LLaMA-Factory 76,000
+nvbn/thefuck 10,267 pandas-dev/pandas 571,849 ytdl-org/youtube-dl 96,157
+TheAlgorithms/ Python 419,917 fastapi/fastapi 177,120 Textualize/rich 45,744
+All-Hands-AI/OpenHands 160,514 Stability-AI/ stablediffusion 13,698 ultralytics/ultralytics 94,201
+gradio-app/gradio 184,722 QuivrHQ/quivr 10,264 freqtrade/freqtrade 188,889
+Appendix F Repo2Run tools
 Showing in Table 13, we design the following actions for Repo2Run to facilitate its invocation.
-Table 6: Command list and their functions
+Table 13: Command list and their functions.
 Command Function
-waitinglist add -p package_name Add item into waiting list. If no “version_constraints” are
-[-v version_constraints] -t tool specified, the latest version will be downloaded by default.
-waitinglist addfile file_path Add all entries from a file similar to requirements.txt format to
-the waiting list. Format should be package_name [version_constraints].
+waitinglist add -p package_name [-v version_constraints] -t tool Add item into waiting list. If no “version_constraints” are specified, the latest version will be downloaded by default.
+waitinglist addfile file_path Add all entries from a file similar to requirements.txt format to the waiting list. Format should be package_name [version_constraints].
 waitinglist clear Clear all items in the waiting list.
-conflictlist solve -v Resolve the conflict for the first element in the conflict list, and
-[version_constraints] update the version constraints for the corresponding package_name and tool to version_constraints.
+conflictlist solve -v[version_constraints] Resolve the conflict for the first element in the conflict list, and update the version constraints for the corresponding package_name and tool to version_constraints.
 If no “version_constraints” are specified, the latest version will be downloaded by default.
@@
 download Download all pending items in the waiting list at once, and the conflict list must be empty before executing.
-runtest Check if the configured environment is correct using“pytest”.
-poetryruntest Check if the configured environment is correct in the poetry environment. If you want to run tests in the poetry environment, run it.
+runtest Check if the built environment is correct using“pytest”.
+poetryruntest Check if the built environment is correct in the poetry environment. If you want to run tests in the poetry environment, run it.
 runpipreqs Generate requirements_pipreqs.txt and pipreqs_output.txt and pipreqs_error.txt.
-change_python_version Switching the Python version in the Docker container will forgo
-python_version any installations made prior to the switch. The Python version number should be represented directly with numbers and dots, without any quotation marks.
-clear_configuration Reset all the configuration to the initial setting of python:3.10.
-Appendix D Benchmark
+change_python_version python_version Switching the Python version in the Docker container will forgo any installations made prior to the switch.
+The Python version number should be represented directly with numbers and dots, without any quotation marks.
+clear_configuration Reset all the buildings to the initial setting of python:3.10.
+Appendix G Repository building success status
+As shown in Table LABEL:tabs:success_status, we list the success status of all the packages in our constructed benchmark.
 Table 14: Success status of each package in the benchmark.
@@
 zipnn/zipnn 007319 Yes zou-group/textgrad b2dc68 Yes
-Appendix E pipreqs baseline settings
+Appendix H Baseline settings
+H.1 Pipreqs settings
 Figure 11 shows the template of a Dockerfile generated using “requirements_pipreqs.txt” created by pipreqs.
 Figure 11: Dockerfile prompt using the “requirements_pipreqs.txt” generated by pipreqs.
+H.2 SWE-agent settings
+Below is the settings of SWE-agent for building the environment:
+agent:
+templates:
+system_template: |-
+SETTING: You are an expert skilled in environment configuration, operating directly in the command line with a specialized interface.
+The specialized interface consists of {{WINDOW}} lines of context at a time. In addition to typical bash commands, you can also use the following commands to help you with environment setup.
+COMMANDS:
+{{command_docs}}
+Please note that THE EDIT COMMAND REQUIRES PROPER INDENTATION. Ensure correct indentation when modifying code.
+RESPONSE FORMAT:
+Your shell prompt is formatted as follows:
+(Current task: <task name>) <cwd> \$
+You need to format your output using two fields: discussion and command.
+Example:
+DISCUSSION
+Ill start by using ls to see the files in the root directory.
+ls /repo
+You should only include a SINGLE command in the command section and then wait for a response. Everything in the DISCUSSION section will be saved for future reference.
+You can use any other bash commands and the special commands listed above, but interactive session commands (e.g., python, vim) are not supported.
+instance_template: |-
+Were currently setting up the environment for the following task. Here are the details:
+TASK:
+{{task_details}}
+INSTRUCTIONS:
+Now, youll carry out this task on your own. Your terminal session has begun in the repositorys root directory. Use the provided commands and any bash commands you need. Edit and check files as needed.
+The goal is to generate a Dockerfile that can successfully build and run the tests in the repository using the command "pytest --collect-only -q".
+NOTE:
+1. The repository is cloned into /repo.
+2. The Dockerfile should start with the following lines (if the base image is python:3.10 and repository is adamobeng/wddbfs):
+FROM python:3.10
+RUN pip install pytest
+RUN git clone https://github.com/adamobeng/wddbfs.git
+RUN mkdir /repo
+RUN git config --global --add safe.directory /repo
+RUN cp -r /wddbfs/. /repo && rm -rf /wddbfs/
+RUN rm -rf /wddbfs
+Your task includes:
+0. **Generate Dockerfile**: Create a file named "Dockerfile" in the root path (e.g, the absolute path is /Dockerfile).
+1. **Read Directory Structure**: Check the folder structure in the root directory.
+2. **Check the Configuration Files**: Inspect files like "requirements.txt", "setup.py", "setup.cfg", "Pipfile*", etc.
+3. **Determine Package Dependencies**: Handle dependencies and manage conflicting dependency versions.
+4. **Testing**: Ensure "pytest /repo --collect-only -q" runs without errors.
+5. **Generate Dockerfile**: Write necessary installation or setup steps determined from the inspection. If you finish run testing successfully, you should modify the Dockerfile in /Dockerfile.
+IMPORTANT TIPS:
+* Check the directory and files carefully.
+* Make sure Dockerfile commands are correct.
+* Use proper Dockerfile syntax and indentation.
+* Test the final Dockerfile by running "pytest /repo --collect-only -q".
+(Current task: Generate Dockerfile)
+(Current directory: {{working_dir}})
+bash-\$
+next_step_template: |-
+{{observation}}
+(Current task: Generate Dockerfile)
+(Current directory: {{working_dir}})
+bash-\$
+next_step_no_output_template: |-
+Your command ran successfully and did not produce any output.
+(Current task: Generate Dockerfile)
+(Current directory: {{working_dir}})
+bash-\$
+tools:
+env_variables:
+WINDOW: 100
+OVERLAP: 2
+bundles:
+- path: tools/registry
+- path: tools/defaults
+- path: tools/search
+- path: tools/edit_linting
+- path: tools/submit
+- path: tools/env_setup
+parse_function:
+type: thought_action
+history_processors:
+- type: last_n_observations
+n: 5
+Appendix I Failure Case
+Figure 12 illustrates a “module not found” error when building the repository “jialuechen/deepfolio”.
+In this case, the issue arises from the absence of updating the unit tests.
+The test file “test_stats.py” attempts to import modules from “deepfolio.stats”, but “stats” does not exist in “deepfolio”.
+Consequently, no matter how the LLM agent operates, it cannot directly run this test.
+This highlights the importance for developers to continuously update existing unit tests as the code repository evolves.
+Figure 12: An example of “module not found” error due to the absence of updating the unit tests in the repository.